### v1

# OM4OV: Leveraging Ontology Matching for Ontology Versioning

Abstract.
Due to the dynamic nature of the semantic web, ontology version control is required to capture time-varying information, most importantly for widely-used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component for efficient ontology management, the growing size of ontologies and accumulating errors caused by manual labour overwhelm current OV approaches. In this paper, we propose yet another approach to performing OV using existing ontology matching (OM) techniques and systems. We introduce a unified OM4OV pipeline. From an OM perspective, we reconstruct a new task formulation, performance measurement, and dataset construction for OV tasks. Reusing the prior alignment(s) from OM, we also propose a cross-reference mechanism to effectively reduce the matching candidature and improve overall OV performance. We experimentally validate the OM4OV pipeline and its cross-reference mechanism using three datasets from the Alignment Evaluation Initiative (OAEI) and exploit insights on OM used for OV tasks.
PVLDB Reference Format:
PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at URL_TO_YOUR_ARTIFACTS.
1. Introduction
Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications (Guarino et al., 2009). An ontology is not static, and the need for version control arises with its birth. While Web data is dynamic, any ontology that is used needs to undergo periodic revisions to keep up with growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation (Klein and Fensel, 2001). For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” (Haller and Polleres, 2020). This may cause undesirable deficiencies in downstream artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues.
The foundation of ontology versioning (OV) aims at distinguishing and recognising changes between different ontology versions. By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology, can apply the correct changes correspondingly (Klein and Fensel, 2001). While various methods for OV have been developed, one approach is to extend the ontology itself with internal version information. An ontology can be issued with a unique identifier (e.g. IRI) or a specific version number (e.g. owl:versionInfo) to be distinguished from other versions; each ontology entity may have a new triple to record its current status (e.g. owl:DeprecatedClass and owl:DeprecatedProperty); or every triple could be extended with a 4th dimension populated with a triple timestamp, similar to the RDF-star schema. Alternatively, version information can be recorded in change logs. Change logs can take the form of notebooks, an extensible markup language (XML), or a knowledge graph (KG). Nevertheless, maintaining version information in the ontology can be time-consuming and labour-intensive. Either extending the current schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or has a separate change log. In such cases, current approaches have limited capabilities in detecting the wrong version information or discovering missing version information.
In this study, we investigate a lightweight and fully automatic version control approach for ontologies. We observe that the nature of OV is very similar to that of ontology matching (OM). Both are introduced to tackle interoperability between ontologies, with ontology entities (mainly classes and properties) serving as inputs. While OM is a well-studied problem (Otero-Cerdeira et al., 2015), a unified pipeline can help extend and reuse existing OM techniques and systems for OV tasks. However, these OM techniques and systems cannot be directly used in OV tasks. OM and OV have some important differences. (1) The input of OM is two distinct ontologies, whereas the input of OV is two different versions of one ontology. The output of OM is a set of entity mappings, whereas the output of OV is two sets: changed entities and unchanged entities. (2) OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity. OM determines matched entities between two different ontologies, while OV determines and distinguishes add, delete, remain, and update entities between different versions in one ontology. (3) In addition, there is a dearth of lab-based datasets for OV evaluation. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none of them contain benchmarks designed for OV tasks.
To address these challenges, our aim is to explore the complementary correlation between these two tasks and to develop a uniform approach to shifting OV tasks towards OM tasks. In this setting, these two tasks become interchangeable and share the same pipeline. Specifically, our key contributions include:
  • We introduce a novel OM4OV pipeline. The pipeline leverages OM for OV with new task formulation, measurement, and a testbed that includes data for benchmarking.
  • Drawing on OM practice, a novel cross-reference mechanism is proposed to optimise the OV candidate selection and improve the overall performance of OM.
  • We implement the OM4OV pipeline as a proof-of-concept system and experimentally evaluate the system in the Ontology Alignment Evaluation Initiative (OAEI) OV testbed.
The remainder of the paper is organised as follows. Section 2 reviews the literature on OV. Section 3 introduces our novel OM4OV pipeline, while Section 4 implements and evaluates the proposed pipeline. We discuss our implications and current limitations in Section 5 and Section 6. Section 7 concludes the paper.
2. Related Work
Version control is recognised as a vital element in ontology management. Different versions of an ontology need to be interoperable so that version changes do not impede the effective and sustainable use of the ontology. There have been two main approaches towards OV that aim to enhance an ontology with the ability to represent different versions and to identify their differences.
One option is to include version information inside the ontology. The Simple HTML Ontology Extensions (SHOE) (Heflin and Hendler, 2000) uses the tag BACKWARD-COMPATIBLE-WITH to record version information. Klein et al. (2002) argue that a carefully-managed version numbering system embedded in the URI of the ontology (and therefore the fully expanded name of entities defined in the ontology) can minimise the impact of adopting updated versions because unchanged entities will be unaffected in practice. These approaches have largely been adopted by the later OWL Web Ontology Language (Bechhofer et al., 2004), where a set of annotation properties related to version information is defined. These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify the entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare the deprecated entities. Later, the ontology language $\tau$OWL (Zekri et al., 2016) was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology. This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple (RDF-star Working Group, 2024).
A second option is to create a separate version log to track version changes. Unlike traditional approaches that use an unstructured plain text file, the authors in (Plessers and De Troyer, 2005) propose a new approach that uses a version ontology with the change definition language (CDL) to create a version log. In (Cardoso et al., 2020), the authors construct an historical knowledge graph (HKG). Storing the version log in the knowledge graph not only avoids repetition, but also enables advanced search functions. The authors in (Sassi et al., 2016) argue that version logs may contain redundancy and inconsistent information. They propose a graph-of-relevance approach for interlinking different version logs and removing less relevant versions.
While current approaches simply record human-generated version information, less attention has been paid to machine-generated version information. In other words, both previous approaches rely on the version information contained in or attached to the ontology. If such information is missing or incorrect, there is no way to automatically detect versioning of ontology concepts. In this study, we introduce a lightweight and fully automatic OV approach. Our approach advances by reusing existing OM systems and techniques for OV tasks, rather than creating a new OV framework from scratch. This paves the way for transferring well-studied OM solutions to OV tasks. With minor modifications required, we can reuse existing OM techniques and systems for OV tasks. To the best of our knowledge, our work is the first attempt to systemically analyse and utilise OM for OV.
3. OM4OV Pipeline
Fig. 1 illustrates the overall OM4OV pipeline. Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), an OM task can be considered as finding an alignment ($A$) that contains a collection of mappings. Similarly, given an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), an OV task can be considered as finding an alignment (A) that contains a collection of mappings over the two different versions. However, while an alignment for OM only considers matched entities, OV tracks both matched and non-matched entities. Further, matched entities are composed of two subsets remain and update, and non-matched entities are composed of add and delete entities. In practice, we expect for OV that unchanged remain entities dominate.
Figure 1. OM4OV pipeline.
3.1. Task Formulation
The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or another relation, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$ (Euzenat, 2007).
Similarly, an OV task can be formalised as finding two variants of an alignment $A_{match}$ and $A_{non-match}$ between ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$. While OM may have different mapping relations, OV narrows down the task and focuses only on the equivalence relation. We classify four subset alignments produced in the OV process, namely $A_{remain}$ ($A\odot$), $A_{update}$ ($A\otimes$), $A_{add}$ ($A\oplus$), and $A_{delete}$ ($A\ominus$).
(1) $A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c\geq s\}=A\odot\cup A\otimes$
$A_{non-match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c<s\}=A\oplus\cup A\ominus$
  1. (1)
    The remain and update entities are actually matched entities between different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions. If the naming convention changes, we will classify ontology entities that have changed names according to systematic convention as update entities, by modelling their confidence $c$ as a high number less than $s$, but close to it. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions. Therefore, $A\odot$ and $A\otimes$ are defined as:
    (2) $A\odot=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c=1\}=A_{match}(c=1)$
    $A\otimes=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c<1\}=A_{match}(c\neq 1)$
  2. (2)
    The add and delete entities are actually non-matched entities between different ontology versions. The add entities are the non-matched terms in $O^{\prime}$, which have no matches in $O$. Similarly, we can interpret the delete entities as non-matched entities in $O$. Therefore, $A\oplus$ and $A\ominus$ are defined as:
    (3) $A\oplus=\{e2\in O^{\prime}$ $\cap A_{non-match}\}=\{e2\in O^{\prime}\}\setminus A_{match}$
    $A\ominus=\{e1\in O$ $\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\}\setminus A_{match}$
3.2. Performance Measurement
Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision, recall, and F1 score. While precision measures matching correctness and recall measures matching completeness, there is an inherent trade-off between precision and recall. The F1 score offers a harmonic mean to balance the matching correctness and completeness. They are defined as:
(4) $Precison=\frac{|A\cap R|}{|A|}\quad Recall=\frac{|A\cap R|}{|R|}$
$F_{1}\ Score=\frac{2}{Precison^{-1}+Recall^{-1}}$
OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance. Within each sub-measure, the precision-recall trade-off still holds. Across different sub-measures, they are not independent but satisfy the following equations, where $N(O)$ is the number of entities in $O$ and $N(O^{\prime})$ is the number of entities in $O^{\prime}$:
(5) $N(O)+N(O^{\prime})$ $=2\times(|A\odot|+|A\otimes|)+|A\oplus|+|A\ominus|$
For each change in a part of an alignment, other parts will change accordingly. For example, if a new alignment is found in $A\odot$ or $A\otimes$, then the number of alignments in $A\oplus$ and $A\ominus$ will be reduced by one each, and vice versa. If we define $\Delta A$ as a universal change in OV, any changes in the alignments satisfy the following equation:
(6) $|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=-|\Delta A\oplus|=-|\Delta A\ominus|$
We define the corresponding changes of $|A\cap R|$ in remain, update, add, and delete as $\Delta(A\cap R)\odot\otimes$ and $\Delta(A\cap R)\oplus\ominus$.
  • For recall, there are direct relations between different sub-measures. Recalldirect-$\odot$ and Recalltensor-$\otimes$ will increase with Recalldirect-$\oplus$ and Recallsymmetric-$\ominus$ decreasing and vice versa.
    (7) $Recall\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}$
    $Recall\oplus\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|R|}$
  • For precision, there are no direct relations between different sub-measures. The indirect relations, described qualitatively in the following as nondecreasing ($\uparrow$), nonincreasing ($\downarrow$), or either (?)?(?)( ? ), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
    (8) $Precision\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$
    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\odot\otimes|\times|A|}<1\quad Precision\odot\otimes\uparrow$
    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\odot\otimes|\times|A|}>1\quad Precision\odot\otimes\downarrow$
    $Precision\oplus\ominus\ ?=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\oplus\ominus|\times|A|}>1\quad Precision\oplus\ominus\uparrow$
    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\oplus\ominus|\times|A|}<1\quad Precision\oplus\ominus\downarrow$
3.3. Dataset Construction
We propose an approach to constructing synthetic OV datasets from OM datasets. Figure 2 illustrates the generation of OM4OV datasets. Our approach is described in the following steps:
Figure 2. Generation of OM4OV datasets.
  1. (1)
    The original OAEI datasets for OM provide two ontologies, $O_{s}$ and $O_{t}$. We choose one to be the intermediate ontology $O_{i}$. We retrieve all ontology entities from $O_{i}$.
  2. (2)
    There are four possible changes in OV tasks: remain, update, add, and delete. Each ontology entity in $O_{i}$ is randomly assigned one of these.
  3. (3)
    For entities assigned to update, we need to generate the updated entity name. We should expect the new name to have a similar meaning to its original name. For example, the entity name “ConferenceVenuePlace” could be replaced with “Conference_hall” or “Conference_building”, but not the general names “Place” or “Location”. To achieve this goal, we retrieve all equivalent entities provided by reference.xml included in the original OAEI datasets and use them as a replacement synonym corpus. For those entities whose names are unique identifiers or codes (and not textually-meaningful names), we use their annotation properties (e.g. rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead. For those entities that do not have synonyms in the generated list, we randomly re-assign the entity to remain, add, or delete.
    For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$. Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$. By construction, we also have that the four sets are pairwise disjoint.
  4. (4)
    Based on the entity assignments, we generate four corresponding versioning references, namely vr-remain.xml, vr-update.xml, vr-add.xml, and vr-delete.xml. For entities assigned to remain, no operation is required. For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list.
  5. (5)
    We generate $O$ and $O^{\prime}$ according to the following two rules:
    1. (a)
      $O$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{add or }p\in\textit{add or }o\in\textit{add}\}$. That is, $O$ is constructed as $O_{i}$ without all the triples related to entities in the add list.
    2. (b)
      Let $e$ be an entity and $A$ be a mapping of the form $\{e_{1}\rightarrow e^{\prime},e_{2}\rightarrow e^{\prime\prime},\ldots e_{n}\rightarrow e^{{}^{\prime\prime}...^{\prime}}\}$. Then $map(e,M)$ is defined to be $e^{\prime}$ if there is an $e^{\prime}$ such that $e\rightarrow e^{\prime}\in A$ and to be $e$ otherwise. Now, $O^{\prime}$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{delete}\textit{ or }p\in\textit{delete}\textit{ or }o\in\textit{delete}\}\cup\{(s^{\prime},p^{\prime},o^{\prime})|(s,p,o)\textit{ is a triple}\in O_{i}$ and $s^{\prime}=map(s,\textit{update})\textit{ or }p^{\prime}=map(p,\textit{update})\textit{ or }o^{\prime}=map(o,\textit{update})\}$. That is, $O^{\prime}$ is constructed as $O_{i}$ without all the triples related to entities in the delete list and updated for all the triples related to entities in the update list.
Unlike the original OAEI datasets for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks.
3.4. Pipeline Optimisation
Often, ontology creators provide cross-references to other ontologies to enhance interoperability. For example, the cross-reference between the CMT ontology and the Conference ontology is provided along with the CMT ontology. Reusing these cross-references developed for OM tasks, we propose a novel mechanism to reduce matching candidates and also to improve overall OV performance.
Figure 3 illustrates the cross-reference mechanism used in the OM4OV pipeline. We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$. This number can be significantly reduced by removing prior matches (i.e. $O\cap O_{r}\cap O^{\prime}$) and non-matches (i.e. $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$). The prior matching will be part of the final alignment, while the known non-matching will be completely removed in the subsequent OV process. The OV process then only determines the posterior alignment. In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities. Matching performance is also improved by using these known mappings. Since $A_{\pi}$ are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain. On the other hand, while the known non-matches are removed from the OV process, this reduces the complexity of detecting the posterior alignment.
Figure 3. Cross-reference mechanism.
Given a reference ontology ($O_{r}$) with an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), two cross-references between $O$ and $O_{r}$ ($R_{or}$) and between $O^{\prime}$ and $O_{r}$ ($R_{o^{\prime}r}$) are defined as:
(9) $R_{or}$ =$=\{(e1,e3,\equiv,c)|e1,e3\in O_{r}\cap O,c\geq s\}$
$R_{o^{\prime}r}$ =$=\{(e2,e3,\equiv,c)|e2,e3\in O_{r}\cap O^{\prime},c\geq s\}$
We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV. We call these mappings a prior alignment ($A_{\pi}$). After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$). Therefore, $A_{match}$ in OV can be decomposed into two parts:
(10) $A_{match}=A_{\pi}\cup A_{\pi^{*}}$
  1. (1)
    $A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$. The equivalence relation is transitive, so for $e1\in O$, $e2\in O^{\prime}$, and $e3\in O_{r}$, if $e1=e3$ and $e2=e3$ then $e1=e2$.
    (11) $A_{\pi}=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},c\geq s\}$
  2. (2)
    $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference. None of these mappings would come from any pairwise intersection of $O$, $O_{r}$, and $O^{\prime}$ because $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$ are pre-defined as non-matched entities, and the matched entities in $O\cap O_{r}\cap O^{\prime}$ have already been captured in the $A(\pi)$. As a result, $A_{\pi^{*}}$ can be defined within a smaller scope:
    (12) $A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},c\geq s\}$
[…]
(13) $A_{\pi}=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup...\cup(R_{orn}\cap R_{o^{\prime}rn})$
$A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}$
$\text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup...O_{rn}$
Our cross-reference mechanism has no impact on our proposed measures for OV. However, our cross-reference mechanism is incorporated into our OV testbed. For this, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules:
  1. (1)
    $R_{or}$ is the original reference.xml removing all the mappings related to add entities.
  2. (2)
    $R_{o^{\prime}r}$ is the original reference.xml removing all the mappings related to delete entities and updating all the mappings related to update entities.
4. Implementation & Evaluation
The OM4OV pipeline is implemented in Agent-OV, a variant of Agent-OM (Qiang et al., 2024). Agent-OM is an agent-powered LLM-based OM system. Its foundation framework is designed for traditional OM tasks. We extend the original framework with the OM4OV pipeline so that it can be used to handle OV tasks.
We use the LLM model gpt-4o-mini for evaluation. As an extension of Agent-OM, Agent-OV also supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT (OpenAI, 2024), Anthropic Claude (Anthropic, 2024), and Mistral AI (Mistral AI, 2024), as well as open-source LLMs Meta Llama (Meta, 2024), Google Gemma (Gemma Team, Google DeepMind, 2024), and Alibaba Qwen (Alibaba, 2024). For the performance of Agent-OV using different LLMs, we refer the reader to (Qiang et al., 2024), where we find that API-accessed LLMs generally perform better than open-source LLMs. The hyperparameter settings are set to $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated from the OV testbed.
4.1. Evaluation of OV testbed
The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks. Table 1 lists the detailed information of the selected OAEI track used for the OV testbed. The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies.
Table 1. Selected OAEI track for the OV testbed.
Track Domain Number of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 3
Figures 45, and 6 show the evaluation of the Agent-OV system on the OV testbed. The results indicate that it is possible to unify the OM and OV tasks. With the necessary modifications, the OM systems can also be used in OV tasks.
  1. (1)
    Our experiments show that the OM system achieves acceptable performance in tracking changes over different versions of the ontologies. Interestingly, the performance of our system on OV tasks is generally better than on OM tasks. This could be due to the fact that the ontologies used in the OV tasks are two versions of the same ontology that share consistent ontology design patterns.
  2. (2)
    For the four micro-level sub-measures, we observe:
    1. (a)
      The performance is highest in remain, followed by add and delete, and it is relatively low in update. This trend is consistent across different tracks and ontologies.
      Figure 4. Evaluation of Anatomy Track.
      Figure 5. Evaluation of Conference Track.
      Figure 6. Evaluation of MSE Track.
    2. (b)
      The measurements for remain are commonly very close to 100% and not statistically significant in OV tasks.
    3. (c)
      The measures for add and delete are generally good. This may be because our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect missing mappings.
    4. (d)
      The measurements for update show scope for improvement. This is because finding non-trivial alignments and appropriate similarity thresholds is not easy. Unlike for OM tasks, we believe mixing remain and update in OV tasks is a common pitfall and can cause skewed performance measurement.
  3. (3)
    For the computational time, we observe a longer computation time for the large-scale ontologies in the Anatomy Track. Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still not sufficient for some OV tasks.
4.2. Evaluation of pipeline optimisation
We apply the cross-reference mechanism on the same alignments that we evaluated in Section 4.1. Figure 7 shows the comparison of the OV performance with and without using the cross-reference (CR) mechanism to detect update entities. We can see that the CR mechanism significantly improves recall and precision in most ontologies, resulting in a solid improvement in the F1 score.
(a) Precision
(b) Recall
(c) F1 Score
Figure 7. Evaluation of cross-reference (CR) mechanism.
4.3. Evaluation of hyperparameter settings
Unlike OM tasks, changes in hyperparameter settings in OV tasks do not lead to a trade-off in precision and recall. Instead, the hyperparameter settings directly influence the sub-measures. Lower similarity thresholds and higher top@k values can result in more update entities being detected, while higher similarity thresholds and lower top@k values may find more add and delete entities. The hyperparameter setting has no impact on remain entities.
5. Discussion
So far, we have experimentally validated the OM4OV pipeline with a novel cross-reference mechanism for pipeline optimisation. Although matching performance has been improved, it still does not achieve 100% as false mappings may still exist. Are these mappings genuinely false?
  1. (1)
    False mappings in OV can be a wrong ontology design choice. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by incorrect name conventions. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially aid in a more effective ontology design.
    cmt:hasAuthor => cmt:writtenBy
    cmt:writtenBy => cmt:isWrittenBy
  2. (2)
    False mappings in OV can be an ambiguous “equivalent” relationship. In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference. However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair. This interpretation is not wrong but follows a different ontology design pattern. It is important to notice that the term ”equivalent” in OV is weaker than that in OM. OV allows for roughly “equivalent” mappings. The entities mapped in OV can slightly alter their meanings in response to changes in the domain knowledge understanding.
    cmt:Chairman => cmt:Chair
    cmt:ConferenceChair => cmt:General_Chair
  3. (3)
    False mappings in OV can be an inappropriate setup on the similarity threshold. In the following example, if $similarity\_threshold=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold is relatively higher in the biomedical domain to ensure each terminology is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
    similarity_threshold = 0.95, cmt:SubjectArea => None
    similarity_threshold = 0.90, cmt:SubjectArea => cmt:Topic
6. Limitations
  1. (1)
    We focus mainly on tracking conceptual changes of classes and properties (e.g. adding, deleting, or updating a class). In practice, there are also internal relationship changes (e.g. changing the domain and range of a class or moving a sibling class to a different parent). These changes are currently only indicated by the changes in the similarity score, while details can only be observed by inspecting the classes and properties. Our future work aims to improve the explainability of changes.
  2. (2)
    While OM has a universal measure, our proposed measures for OV have four sub-measures. It is necessary to find a single universal measure combining these sub-measures so that we can fairly assess and compare the system performance in OV and OM. We plan to investigate a merged formula for OV sub-measures in the future. For example, using a harmonic mean to combine sub-measures.
7. Conclusion
In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with the necessary modifications. We propose a novel OM4OV pipeline with a cross-reference mechanism that leverages OM for OV. The new pipeline (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the matching candidates, and (3) improves overall performance. We incorporate the OM4OV pipeline into a new OV system called Agent-OV, a functional extension of Agent-OM to handle OV tasks. Evaluations of three OAEI datasets validate the feasibility and reliability of our system. We also argue that the false mappings detected by OV systems are not necessarily actual false mappings.
Our approach is compatible with ontologies using different versioning methods (using URIs or additional versioning triples), and even ontologies missing or without version information. Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies.
Acknowledgements.
The authors thank Weiqing Wang from Monash University for giving helpful advice on this work. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v2

@@
 OM4OV: Leveraging Ontology Matching for Ontology Versioning
-Abstract.
+Abstract
 Due to the dynamic nature of the semantic web, ontology version control is required to capture time-varying information, most importantly for widely-used ontologies.
@@
 We introduce a unified OM4OV pipeline.
-From an OM perspective, we reconstruct a new task formulation, performance measurement, and dataset construction for OV tasks.
-Reusing the prior alignment(s) from OM, we also propose a cross-reference mechanism to effectively reduce the matching candidature and improve overall OV performance.
-We experimentally validate the OM4OV pipeline and its cross-reference mechanism using three datasets from the Alignment Evaluation Initiative (OAEI) and exploit insights on OM used for OV tasks.
-PVLDB Reference Format:
-PVLDB, 18(1): XXX-XXX, 2025.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at URL_TO_YOUR_ARTIFACTS.
-1. Introduction
-Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications (Guarino et al., 2009).
+From an OM perspective, we reconstruct a new task formulation, measurement, and testbed for OV tasks.
+Reusing the prior alignment(s) from OM, we propose a pipeline optimisation method called cross-reference (CR) mechanism to improve overall OV performance.
+We experimentally validate the OM4OV pipeline and the cross-reference mechanism in modified Ontology Alignment Evaluation Initiative (OAEI) datasets.
+We also discuss the insights on OM used for OV tasks, where some false mappings detected by OV systems are not actually false.
+Index Terms:
+ontology matching, ontology versioning
+I Introduction
+Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications [1].
 An ontology is not static, and the need for version control arises with its birth.
-While Web data is dynamic, any ontology that is used needs to undergo periodic revisions to keep up with growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation (Klein and Fensel, 2001).
-For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” (Haller and Polleres, 2020).
-This may cause undesirable deficiencies in downstream artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues.
+While Web data is dynamic, any ontology that is used needs to undergo periodic revisions to keep up with growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation [2].
+For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [3].
+This may cause undesirable deficiencies in artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
 The foundation of ontology versioning (OV) aims at distinguishing and recognising changes between different ontology versions.
-By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology, can apply the correct changes correspondingly (Klein and Fensel, 2001).
+By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology, can apply the correct changes correspondingly [2].
 While various methods for OV have been developed, one approach is to extend the ontology itself with internal version information.
@@
 Both are introduced to tackle interoperability between ontologies, with ontology entities (mainly classes and properties) serving as inputs.
-While OM is a well-studied problem (Otero-Cerdeira et al., 2015), a unified pipeline can help extend and reuse existing OM techniques and systems for OV tasks.
-However, these OM techniques and systems cannot be directly used in OV tasks.
-OM and OV have some important differences.
-(1) The input of OM is two distinct ontologies, whereas the input of OV is two different versions of one ontology.
+While OM is a well-studied problem [4], a unified pipeline that is able to extend and reuse existing OM techniques and systems for OV tasks would be very desirable.
+However, OM techniques and systems cannot be directly used in OV tasks because OM and OV have some important differences.
+(1) The input of OM is two distinct ontologies, whereas the input of OV is two different versions of one single ontology.
 The output of OM is a set of entity mappings, whereas the output of OV is two sets: changed entities and unchanged entities.
@@
 OM determines matched entities between two different ontologies, while OV determines and distinguishes add, delete, remain, and update entities between different versions in one ontology.
-(3) In addition, there is a dearth of lab-based datasets for OV evaluation.
-The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none of them contain benchmarks designed for OV tasks.
 To address these challenges, our aim is to explore the complementary correlation between these two tasks and to develop a uniform approach to shifting OV tasks towards OM tasks.
-In this setting, these two tasks become interchangeable and share the same pipeline.
 Specifically, our key contributions include:
-  • We introduce a novel OM4OV pipeline. The pipeline leverages OM for OV with new task formulation, measurement, and a testbed that includes data for benchmarking.
-  • Drawing on OM practice, a novel cross-reference mechanism is proposed to optimise the OV candidate selection and improve the overall performance of OM.
-  • We implement the OM4OV pipeline as a proof-of-concept system and experimentally evaluate the system in the Ontology Alignment Evaluation Initiative (OAEI) OV testbed.
+  • We systematically analyse OV from an OM perspective and introduce a novel OM4OV pipeline.
+The pipeline leverages OM for OV with new task formulation, measurement, and a testbed that includes data for benchmarking.
+  • Drawing on OM practice, a novel cross-reference mechanism is proposed to optimise candidate selection and improve the overall performance of OV.
+  • We implement the OM4OV pipeline as a proof-of-concept system and experimentally evaluate its performance.
 The remainder of the paper is organised as follows.
-Section 2 reviews the literature on OV.
-Section 3 introduces our novel OM4OV pipeline, while Section 4 implements and evaluates the proposed pipeline.
-We discuss our implications and current limitations in Section 5 and Section 6.
-Section 7 concludes the paper.
-2. Related Work
+Section II reviews the literature on OV.
+Section III introduces our novel OM4OV pipeline, while Section IV implements and evaluates the proposed pipeline.
+We discuss our implications and current limitations in Section V and Section VI.
+Section VII concludes the paper.
+II Related Work
 Version control is recognised as a vital element in ontology management.
@@
 One option is to include version information inside the ontology.
-The Simple HTML Ontology Extensions (SHOE) (Heflin and Hendler, 2000) uses the tag BACKWARD-COMPATIBLE-WITH to record version information.
-Klein et al.
-(2002) argue that a carefully-managed version numbering system embedded in the URI of the ontology (and therefore the fully expanded name of entities defined in the ontology) can minimise the impact of adopting updated versions because unchanged entities will be unaffected in practice.
-These approaches have largely been adopted by the later OWL Web Ontology Language (Bechhofer et al., 2004), where a set of annotation properties related to version information is defined.
+The Simple HTML Ontology Extensions (SHOE) [5] uses the tag BACKWARD-COMPATIBLE-WITH to record version information.
+The authors in [6] argue that a carefully-managed version numbering system embedded in the URI of the ontology (and therefore the fully expanded name of entities defined in the ontology) can minimise the impact of adopting updated versions because unchanged entities will be unaffected in practice.
+These approaches have largely been adopted by the later OWL Web Ontology Language [7], where a set of annotation properties related to version information is defined.
 These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify the entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare the deprecated entities.
-Later, the ontology language $\tau$OWL (Zekri et al., 2016) was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology.
-This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple (RDF-star Working Group, 2024).
+Later, the ontology language $\tau$OWL [8] was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology.
+This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple [9].
 A second option is to create a separate version log to track version changes.
-Unlike traditional approaches that use an unstructured plain text file, the authors in (Plessers and De Troyer, 2005) propose a new approach that uses a version ontology with the change definition language (CDL) to create a version log.
-In (Cardoso et al., 2020), the authors construct an historical knowledge graph (HKG).
+Unlike traditional approaches that use an unstructured plain text file, the authors in [10] propose a new approach that uses a version ontology with the change definition language (CDL) to create a version log.
+In [11], the authors construct an historical knowledge graph (HKG).
 Storing the version log in the knowledge graph not only avoids repetition, but also enables advanced search functions.
-The authors in (Sassi et al., 2016) argue that version logs may contain redundancy and inconsistent information.
+The authors in [12] argue that version logs may contain redundancy and inconsistent information.
 They propose a graph-of-relevance approach for interlinking different version logs and removing less relevant versions.
@@
 With minor modifications required, we can reuse existing OM techniques and systems for OV tasks.
-To the best of our knowledge, our work is the first attempt to systemically analyse and utilise OM for OV.
-3. OM4OV Pipeline
+To the best of our knowledge, our work is the first to systemically analyse and utilise OM for OV.
+III OM4OV Pipeline
 Fig.
-1 illustrates the overall OM4OV pipeline.
+1 illustrates the overview of OM4OV pipeline.
 Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), an OM task can be considered as finding an alignment ($A$) that contains a collection of mappings.
@@
 In practice, we expect for OV that unchanged remain entities dominate.
-Figure 1. OM4OV pipeline.
-3.1. Task Formulation
-The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or another relation, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$ (Euzenat, 2007).
+Figure 1: OM4OV pipeline.
+III-A Task Formulation
+The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or other relations, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$ [13].
+$\footnotesize A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$ (1)
 Similarly, an OV task can be formalised as finding two variants of an alignment $A_{match}$ and $A_{non-match}$ between ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$.
@@
 We classify four subset alignments produced in the OV process, namely $A_{remain}$ ($A\odot$), $A_{update}$ ($A\otimes$), $A_{add}$ ($A\oplus$), and $A_{delete}$ ($A\ominus$).
-(1) $A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c\geq s\}=A\odot\cup A\otimes$
+$A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c\geq s\}=A\odot\cup A\otimes$ (2)
 $A_{non-match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c<s\}=A\oplus\cup A\ominus$
-  1. (1)
+  1. 1.
     The remain and update entities are actually matched entities between different ontology versions.
 The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions.
-If the naming convention changes, we will classify ontology entities that have changed names according to systematic convention as update entities, by modelling their confidence $c$ as a high number less than $s$, but close to it.
+If the naming convention changes, we will classify ontology entities that have changed names according to systematic convention as update entities, by modelling their confidence $c$ as a high number more than $s$ but less than 1.
 The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions.
 Therefore, $A\odot$ and $A\otimes$ are defined as:
-    (2) $A\odot=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c=1\}=A_{match}(c=1)$
-    $A\otimes=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c<1\}=A_{match}(c\neq 1)$
-  2. (2)
+    $A\odot=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c=1\}=A_{match}(c=1)$ (3)
+    $A\otimes=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},s\leq c<1\}=A_{match}(c\neq 1)$
+  2. 2.
     The add and delete entities are actually non-matched entities between different ontology versions.
@@
 Therefore, $A\oplus$ and $A\ominus$ are defined as:
-    (3) $A\oplus=\{e2\in O^{\prime}$ $\cap A_{non-match}\}=\{e2\in O^{\prime}\}\setminus A_{match}$
-    $A\ominus=\{e1\in O$ $\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\}\setminus A_{match}$
-3.2. Performance Measurement
+    $A\oplus=\{e2\in O^{\prime}$ $\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$ (4)
+    $A\ominus=\{e1\in O$ $\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
+III-B Measurement
 Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision, recall, and F1 score.
 While precision measures matching correctness and recall measures matching completeness, there is an inherent trade-off between precision and recall.
-The F1 score offers a harmonic mean to balance the matching correctness and completeness.
+The F1 score offers a harmonic mean to balance the correctness and completeness.
 They are defined as:
-(4) $Precison=\frac{|A\cap R|}{|A|}\quad Recall=\frac{|A\cap R|}{|R|}$
+$Precison=\frac{|A\cap R|}{|A|}\quad Recall=\frac{|A\cap R|}{|R|}$ (5)
 $F_{1}\ Score=\frac{2}{Precison^{-1}+Recall^{-1}}$
@@
 Across different sub-measures, they are not independent but satisfy the following equations, where $N(O)$ is the number of entities in $O$ and $N(O^{\prime})$ is the number of entities in $O^{\prime}$:
-(5) $N(O)+N(O^{\prime})$ $=2\times(|A\odot|+|A\otimes|)+|A\oplus|+|A\ominus|$
+$N(O)+N(O^{\prime})$ $=2\times(|A\odot|+|A\otimes|)+|A\oplus|+|A\ominus|$ (6)
 For each change in a part of an alignment, other parts will change accordingly.
@@
 If we define $\Delta A$ as a universal change in OV, any changes in the alignments satisfy the following equation:
-(6) $|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=-|\Delta A\oplus|=-|\Delta A\ominus|$
+$|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=-|\Delta A\oplus|=-|\Delta A\ominus|$ (7)
 We define the corresponding changes of $|A\cap R|$ in remain, update, add, and delete as $\Delta(A\cap R)\odot\otimes$ and $\Delta(A\cap R)\oplus\ominus$.
@@
 Recalldirect-$\odot$ and Recalltensor-$\otimes$ will increase with Recalldirect-$\oplus$ and Recallsymmetric-$\ominus$ decreasing and vice versa.
-    (7) $Recall\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}$
+    $Recall\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}$ (8)
     $Recall\oplus\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|R|}$
@@
 ), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
-    (8) $Precision\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$
+    $Precision\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (9)
     $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\odot\otimes|\times|A|}<1\quad Precision\odot\otimes\uparrow$
@@
     $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\oplus\ominus|\times|A|}<1\quad Precision\oplus\ominus\downarrow$
-3.3. Dataset Construction
-We propose an approach to constructing synthetic OV datasets from OM datasets. Figure 2 illustrates the generation of OM4OV datasets. Our approach is described in the following steps:
-Figure 2. Generation of OM4OV datasets.
-  1. (1)
+III-C Testbed
+There is a dearth of lab-based datasets for OV evaluation.
+The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none of them contain benchmarks designed for OV tasks.
+We propose an approach to constructing synthetic OV datasets from OM datasets.
+Fig.
+2 illustrates the generation of OM4OV datasets.
+Our approach is described in the following steps:
+Figure 2: Generation of OM4OV datasets.
+  1. 1.
     The original OAEI datasets for OM provide two ontologies, $O_{s}$ and $O_{t}$. We choose one to be the intermediate ontology $O_{i}$. We retrieve all ontology entities from $O_{i}$.
-  2. (2)
-    There are four possible changes in OV tasks: remain, update, add, and delete. Each ontology entity in $O_{i}$ is randomly assigned one of these.
-  3. (3)
+  2. 2.
+    There are four possible changes in OV tasks: remain, update, add, and delete.
+Each ontology entity in $O_{i}$ is randomly assigned one of these.
+For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$.
+Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$.
+By construction, we also have the four sets as pairwise disjoints.
+  3. 3.
     For entities assigned to update, we need to generate the updated entity name.
@@
 For those entities that do not have synonyms in the generated list, we randomly re-assign the entity to remain, add, or delete.
-    For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$.
-Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$.
-By construction, we also have that the four sets are pairwise disjoint.
-  4. (4)
+  4. 4.
     Based on the entity assignments, we generate four corresponding versioning references, namely vr-remain.xml, vr-update.xml, vr-add.xml, and vr-delete.xml.
@@
 For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list.
-  5. (5)
+  5. 5.
     We generate $O$ and $O^{\prime}$ according to the following two rules:
@@
 That is, $O^{\prime}$ is constructed as $O_{i}$ without all the triples related to entities in the delete list and updated for all the triples related to entities in the update list.
+  6. 6.
+    We also incorporate the cross-reference mechanism into our OV testbed, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules: (1) $R_{or}$ is the original reference.xml that removes all the mappings related to add entities.
+(2) $R_{o^{\prime}r}$ is the original reference.xml that removes all the mappings related to delete entities and updates all the mappings related to update entities.
 Unlike the original OAEI datasets for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed.
@@
 For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks.
-3.4. Pipeline Optimisation
+The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks.
+Table I lists the detailed information of the selected OAEI track used for the OV testbed.
+The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies.
+TABLE I: Selected OAEI track for the OV testbed.
+Track Domain Number of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 3
+III-D Cross-Reference Mechanism
 Often, ontology creators provide cross-references to other ontologies to enhance interoperability.
@@
 Reusing these cross-references developed for OM tasks, we propose a novel mechanism to reduce matching candidates and also to improve overall OV performance.
-Figure 3 illustrates the cross-reference mechanism used in the OM4OV pipeline.
+Fig.
+3 illustrates the cross-reference mechanism used in the OM4OV pipeline.
 We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$.
@@
 Matching performance is also improved by using these known mappings.
-Since $A_{\pi}$ are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain.
+Since prior matches are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain.
 On the other hand, while the known non-matches are removed from the OV process, this reduces the complexity of detecting the posterior alignment.
-Figure 3. Cross-reference mechanism.
+Figure 3: Cross-reference mechanism.
 Given a reference ontology ($O_{r}$) with an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), two cross-references between $O$ and $O_{r}$ ($R_{or}$) and between $O^{\prime}$ and $O_{r}$ ($R_{o^{\prime}r}$) are defined as:
-(9) $R_{or}$ =$=\{(e1,e3,\equiv,c)|e1,e3\in O_{r}\cap O,c\geq s\}$
+$R_{or}$ =$=\{(e1,e3,\equiv,c)|e1,e3\in O_{r}\cap O,c\geq s\}$ (10)
 $R_{o^{\prime}r}$ =$=\{(e2,e3,\equiv,c)|e2,e3\in O_{r}\cap O^{\prime},c\geq s\}$
@@
 Therefore, $A_{match}$ in OV can be decomposed into two parts:
-(10) $A_{match}=A_{\pi}\cup A_{\pi^{*}}$
-  1. (1)
+$A_{match}=A_{\pi}\cup A_{\pi^{*}}$ (11)
+  1. 1.
     $A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$.
 The equivalence relation is transitive, so for $e1\in O$, $e2\in O^{\prime}$, and $e3\in O_{r}$, if $e1=e3$ and $e2=e3$ then $e1=e2$.
-    (11) $A_{\pi}=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},c\geq s\}$
-  2. (2)
+    $A_{\pi}=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},c\geq s\}$ (12)
+  2. 2.
     $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference.
@@
 As a result, $A_{\pi^{*}}$ can be defined within a smaller scope:
-    (12) $A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},c\geq s\}$
+    $A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},c\geq s\}$ (13)
 […]
-(13) $A_{\pi}=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup...\cup(R_{orn}\cap R_{o^{\prime}rn})$
+$A_{\pi}=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup...\cup(R_{orn}\cap R_{o^{\prime}rn})$ (14)
 $A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}$
 $\text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup...O_{rn}$
-Our cross-reference mechanism has no impact on our proposed measures for OV.
-However, our cross-reference mechanism is incorporated into our OV testbed.
-For this, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules:
-  1. (1)
-    $R_{or}$ is the original reference.xml removing all the mappings related to add entities.
-  2. (2)
-    $R_{o^{\prime}r}$ is the original reference.xml removing all the mappings related to delete entities and updating all the mappings related to update entities.
-4. Implementation & Evaluation
-The OM4OV pipeline is implemented in Agent-OV, a variant of Agent-OM (Qiang et al., 2024).
+IV Implementation & Evaluation
+The OM4OV pipeline is implemented in Agent-OV, a variant of Agent-OM [14].
 Agent-OM is an agent-powered LLM-based OM system.
@@
 We extend the original framework with the OM4OV pipeline so that it can be used to handle OV tasks.
-We use the LLM model gpt-4o-mini for evaluation.
-As an extension of Agent-OM, Agent-OV also supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT (OpenAI, 2024), Anthropic Claude (Anthropic, 2024), and Mistral AI (Mistral AI, 2024), as well as open-source LLMs Meta Llama (Meta, 2024), Google Gemma (Gemma Team, Google DeepMind, 2024), and Alibaba Qwen (Alibaba, 2024).
-For the performance of Agent-OV using different LLMs, we refer the reader to (Qiang et al., 2024), where we find that API-accessed LLMs generally perform better than open-source LLMs.
-The hyperparameter settings are set to $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated from the OV testbed.
-4.1. Evaluation of OV testbed
-The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks.
-Table 1 lists the detailed information of the selected OAEI track used for the OV testbed.
-The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies.
-Table 1. Selected OAEI track for the OV testbed.
-Track Domain Number of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 3
-Figures 45, and 6 show the evaluation of the Agent-OV system on the OV testbed.
+As an extension of Agent-OM, Agent-OV also supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [15], Anthropic Claude [16], and Mistral AI [17], as well as open-source LLMs Meta Llama [18], Google Gemma [19], and Alibaba Qwen [20].
+For the performance of Agent-OV using different LLMs, we refer the reader to [14], where we find that API-accessed LLMs generally perform better than open-source LLMs.
+IV-A Evaluation of OV testbed
+Fig.
+4 shows the evaluation of the Agent-OV system on the OV testbed.
+We use the OpenAI model gpt-4o-mini (2024-07-18) [21] for evaluation.
+The embedding chosen correspondingly is OpenAI text-embedding-ada-002 [22].
+The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated.
+We run the OV testbed for 5 iterations, simulating 5 distinct versions for each ontology.
 The results indicate that it is possible to unify the OM and OV tasks.
 With the necessary modifications, the OM systems can also be used in OV tasks.
-  1. (1)
+(a) Anatomy Track: human
+(b) Anatomy Track: mouse
+(c) Conference Track: cmt
+(d) Conference Track: conference
+(e) Conference Track: confof
+(f) Conference Track: edas
+(g) Conference Track: ekaw
+(h) Conference Track: iasted
+(i) Conference Track: sigkdd
+(j) MSE Track: EMMO
+(k) MSE Track: MatInfo
+(l) MSE Track: MatOnto
+Figure 4: Evaluation of OV testbed.
+  1. 1.
     Our experiments show that the OM system achieves acceptable performance in tracking changes over different versions of the ontologies.
@@
 This could be due to the fact that the ontologies used in the OV tasks are two versions of the same ontology that share consistent ontology design patterns.
-  2. (2)
+  2. 2.
     For the four micro-level sub-measures, we observe:
@@
       The performance is highest in remain, followed by add and delete, and it is relatively low in update. This trend is consistent across different tracks and ontologies.
-      Figure 4. Evaluation of Anatomy Track.
-      Figure 5. Evaluation of Conference Track.
-      Figure 6. Evaluation of MSE Track.
     2. (b)
@@
 Unlike for OM tasks, we believe mixing remain and update in OV tasks is a common pitfall and can cause skewed performance measurement.
-  3. (3)
+  3. 3.
     For the computational time, we observe a longer computation time for the large-scale ontologies in the Anatomy Track.
 Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still not sufficient for some OV tasks.
-4.2. Evaluation of pipeline optimisation
-We apply the cross-reference mechanism on the same alignments that we evaluated in Section 4.1.
-Figure 7 shows the comparison of the OV performance with and without using the cross-reference (CR) mechanism to detect update entities.
-We can see that the CR mechanism significantly improves recall and precision in most ontologies, resulting in a solid improvement in the F1 score.
-(a) Precision
-(b) Recall
-(c) F1 Score
-Figure 7. Evaluation of cross-reference (CR) mechanism.
-4.3. Evaluation of hyperparameter settings
-Unlike OM tasks, changes in hyperparameter settings in OV tasks do not lead to a trade-off in precision and recall.
-Instead, the hyperparameter settings directly influence the sub-measures.
-Lower similarity thresholds and higher top@k values can result in more update entities being detected, while higher similarity thresholds and lower top@k values may find more add and delete entities.
-The hyperparameter setting has no impact on remain entities.
-5. Discussion
+IV-B Evaluation of cross-reference mechanism
+Fig.
+5 shows the comparison of the OV performance with and without using the cross-reference (CR) mechanism on the same OV testbed we analysed in Section IV-A.
+We can see that the CR mechanism significantly improves the capability of Agent-OV to detect update entities and slightly improves the performance on detecting add and delete entities, with no effect on the remain entities.
+More specifically, the performance improvement is mainly caused by the recall for update entities, while this is more evident in the precision for add and delete entities.
+(a) Remain-Precision
+(b) Update-Precision
+(c) Add-Precision
+(d) Delete-Precision
+(e) Remain-Recall
+(f) Update-Recall
+(g) Add-Recall
+(h) Delete-Recall
+(i) Remain-F1 Score
+(j) Update-F1 Score
+(k) Add-F1 Score
+(l) Delete-F1 Score
+Figure 5: Evaluation of cross-reference (CR) mechanism.
+IV-C Evaluation of hyperparameter settings
+Fig.
+6 shows the comparison of using different similarity thresholds in the CMT Ontology versioning.
+Similar to OM tasks, changes in hyperparameter settings in OV tasks will lead to a trade-off in precision and recall.
+Moreover, the hyperparameter settings also influence the sub-measures.
+We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities.
+Also, detecting update entities and detecting add and delete is negatively correlated.
+Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities.
+The result of the experiment also shows that our cross-reference mechanism is less sensitive to the similarity threshold, which means that it can be useful in scenarios where the optimal similarity threshold is unclear or hard to determine.
+Figure 6: Evaluation of different similarity thresholds.
+V Discussion
 So far, we have experimentally validated the OM4OV pipeline with a novel cross-reference mechanism for pipeline optimisation.
 Although matching performance has been improved, it still does not achieve 100% as false mappings may still exist.
-Are these mappings genuinely false?
-  1. (1)
+Are these mappings actually false?
+  1. 1.
     False mappings in OV can be a wrong ontology design choice.
@@
 Ideally, we should avoid this type of ontology design.
-This example also demonstrates a unique benefit of using OM for OV, which could potentially aid in a more effective ontology design.
+This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
     cmt:hasAuthor => cmt:writtenBy
     cmt:writtenBy => cmt:isWrittenBy
-  2. (2)
+  2. 2.
     False mappings in OV can be an ambiguous “equivalent” relationship.
@@
     cmt:ConferenceChair => cmt:General_Chair
-  3. (3)
+  3. 3.
     False mappings in OV can be an inappropriate setup on the similarity threshold.
-In the following example, if $similarity\_threshold=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$.
+In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$.
 Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent.
 For example, the similarity threshold is relatively higher in the biomedical domain to ensure each terminology is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
-    similarity_threshold = 0.95, cmt:SubjectArea => None
-    similarity_threshold = 0.90, cmt:SubjectArea => cmt:Topic
-6. Limitations
-  1. (1)
-    We focus mainly on tracking conceptual changes of classes and properties (e.g.
-adding, deleting, or updating a class).
+    s = 0.95, cmt:SubjectArea => None
+    s = 0.90, cmt:SubjectArea => cmt:Topic
+VI Limitations
+We focus mainly on tracking conceptual changes in classes and properties (e.g.
+adding, deleting, or updating).
 In practice, there are also internal relationship changes (e.g.
@@
 Our future work aims to improve the explainability of changes.
-  2. (2)
 While OM has a universal measure, our proposed measures for OV have four sub-measures.
@@
 For example, using a harmonic mean to combine sub-measures.
-7. Conclusion
+VII Conclusion
 In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with the necessary modifications.
@@
 Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies.
-Acknowledgements.
-The authors thank Weiqing Wang from Monash University for giving helpful advice on this work.
+In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time (e.g.
+smart building ontologies Brick Schema [23], RealEstateCore [24], and ASHRAE Standard 223P [25]).
+Supplemental Material Statement
+Agent-OV is an extended version of Agent-OM.
+Agent-OM can be downloaded from https://github.com/qzc438-research/ontology-llm.
+The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01).
+The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
+Acknowledgment
+The authors would like to thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper.
 The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v3

@@
-OM4OV: Leveraging Ontology Matching for Ontology Versioning
+
 Abstract
-Due to the dynamic nature of the semantic web, ontology version control is required to capture time-varying information, most importantly for widely-used ontologies.
+Due to the dynamic nature of the Semantic Web, version control is necessary to capture time-varying information, particularly for widely used ontologies.
 Despite the long-standing recognition of ontology versioning (OV) as a crucial component for efficient ontology management, the growing size of ontologies and accumulating errors caused by manual labour overwhelm current OV approaches.
@@
 We introduce a unified OM4OV pipeline.
-From an OM perspective, we reconstruct a new task formulation, measurement, and testbed for OV tasks.
-Reusing the prior alignment(s) from OM, we propose a pipeline optimisation method called cross-reference (CR) mechanism to improve overall OV performance.
-We experimentally validate the OM4OV pipeline and the cross-reference mechanism in modified Ontology Alignment Evaluation Initiative (OAEI) datasets.
-We also discuss the insights on OM used for OV tasks, where some false mappings detected by OV systems are not actually false.
-Index Terms:
+From an OM perspective, we reconstruct a new task formulation and measurement for OV tasks.
+Building upon the prior alignment(s) from OM, we propose a pipeline optimisation method called the cross-reference (CR) mechanism to enhance overall OV performance.
+We experimentally validate the OM4OV pipeline and the cross-reference mechanism in the OV tested originating from the Ontology Alignment Evaluation Initiative (OAEI) datasets.
+We also discuss insights into OM used for OV tasks, where some false mappings detected by OV systems are not actually untrue.
+keywords:
 ontology matching, ontology versioning
-I Introduction
+category:
+[…]
+1 Introduction
 Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications [9].
 An ontology is not static, and the need for version control arises with its birth.
-While Web data is dynamic, any ontology that is used needs to undergo periodic revisions to keep up with growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation [2].
+While web data is dynamic, any ontology used must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation [14].
 For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [10].
 This may cause undesirable deficiencies in artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
-The foundation of ontology versioning (OV) aims at distinguishing and recognising changes between different ontology versions.
+Ontology versioning (OV) aims to distinguish and recognise changes between different ontology versions.
 By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology, can apply the correct changes correspondingly [14].
@@
 Change logs can take the form of notebooks, an extensible markup language (XML), or a knowledge graph (KG).
-Nevertheless, maintaining version information in the ontology can be time-consuming and labour-intensive.
+However, maintaining version information in the ontology can be time-consuming and labour-intensive.
 Either extending the current schema or using change logs requires consistent updating over time.
@@
 Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or has a separate change log.
-In such cases, current approaches have limited capabilities in detecting the wrong version information or discovering missing version information.
+In such cases, current approaches have limited capabilities in detecting incorrect version information or identifying missing version information.
 In this study, we investigate a lightweight and fully automatic version control approach for ontologies.
 We observe that the nature of OV is very similar to that of ontology matching (OM).
-Both are introduced to tackle interoperability between ontologies, with ontology entities (mainly classes and properties) serving as inputs.
-While OM is a well-studied problem [4], a unified pipeline that is able to extend and reuse existing OM techniques and systems for OV tasks would be very desirable.
-However, OM techniques and systems cannot be directly used in OV tasks because OM and OV have some important differences.
-(1) The input of OM is two distinct ontologies, whereas the input of OV is two different versions of one single ontology.
+Both are introduced to address interoperability between ontologies, with ontology entities serving as inputs.
+While OM is a well-studied problem [19], a unified pipeline that can extend and reuse existing OM techniques and systems for OV tasks would be highly desirable.
+However, OM techniques and systems cannot be directly applied to OV tasks because OM and OV have some significant differences.
+(1) The OM input is two distinct ontologies, while the OV input is two different versions of a single ontology.
 The output of OM is a set of entity mappings, whereas the output of OV is two sets: changed entities and unchanged entities.
 (2) OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity.
-OM determines matched entities between two different ontologies, while OV determines and distinguishes add, delete, remain, and update entities between different versions in one ontology.
-To address these challenges, our aim is to explore the complementary correlation between these two tasks and to develop a uniform approach to shifting OV tasks towards OM tasks.
+OM determines matched entities between two different ontologies, while OV distinguishes add, delete, remain, and update entities between different versions in one ontology.
+To address these challenges, our goal is to investigate the complementary relationship between these two tasks and to develop a unified approach to shifting OV tasks towards OM tasks.
 Specifically, our key contributions include:
   • We systematically analyse OV from an OM perspective and introduce a novel OM4OV pipeline.
-The pipeline leverages OM for OV with new task formulation, measurement, and a testbed that includes data for benchmarking.
-  • Drawing on OM practice, a novel cross-reference mechanism is proposed to optimise candidate selection and improve the overall performance of OV.
-  • We implement the OM4OV pipeline as a proof-of-concept system and experimentally evaluate its performance.
+  • Drawing on OM practice, a novel cross-reference mechanism is proposed to optimise candidate selection and improve the overall performance of the OM4OV pipeline.
+  • We implement the OM4OV pipeline and the cross-reference mechanism in a proof-of-concept system and experimentally evaluate its performance.
+  • We find that some false mappings detected by OV systems are not actually untrue, but can be caused by a flawed choice of ontology design, an ambiguous “equivalent” relationship, or an inappropriate setting of the similarity threshold.
 The remainder of the paper is organised as follows.
-Section II reviews the literature on OV.
-Section III introduces our novel OM4OV pipeline, while Section IV implements and evaluates the proposed pipeline.
-We discuss our implications and current limitations in Section V and Section VI.
-Section VII concludes the paper.
-II Related Work
+Section 2 reviews the literature on OV.
+Section 3 introduces our novel OM4OV pipeline, while Section 4 evaluates the proposed pipeline.
+We propose our pipeline optimisation method (i.e.
+cross-reference mechanism) and experimentally evaluate its performance in Section 5.
+Section 6 and 7 discuss our implications and current limitations.
+Section 8 concludes the paper.
+2 Related Work
 Version control is recognised as a vital element in ontology management.
@@
 There have been two main approaches towards OV that aim to enhance an ontology with the ability to represent different versions and to identify their differences.
-One option is to include version information inside the ontology.
+One option is to include version information within the ontology.
 The Simple HTML Ontology Extensions (SHOE) [12] uses the tag BACKWARD-COMPATIBLE-WITH to record version information.
@@
 This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple [22].
-A second option is to create a separate version log to track version changes.
+A second option is to create a separate version log to track changes in versions.
 Unlike traditional approaches that use an unstructured plain text file, the authors in [20] propose a new approach that uses a version ontology with the change definition language (CDL) to create a version log.
-In [11], the authors construct an historical knowledge graph (HKG).
+In [5], the authors construct a historical knowledge graph (HKG).
 Storing the version log in the knowledge graph not only avoids repetition, but also enables advanced search functions.
-The authors in [12] argue that version logs may contain redundancy and inconsistent information.
+The authors in [23] argue that the version logs may contain redundancy and inconsistent information.
 They propose a graph-of-relevance approach for interlinking different version logs and removing less relevant versions.
@@
 Our approach advances by reusing existing OM systems and techniques for OV tasks, rather than creating a new OV framework from scratch.
-This paves the way for transferring well-studied OM solutions to OV tasks.
-With minor modifications required, we can reuse existing OM techniques and systems for OV tasks.
-To the best of our knowledge, our work is the first to systemically analyse and utilise OM for OV.
-III OM4OV Pipeline
-Fig.
-1 illustrates the overview of OM4OV pipeline.
+To the best of our knowledge, our work is the first to systematically analyse and utilise OM for OV.
+3 OM4OV Pipeline
+Figure 1 illustrates the overview of the OM4OV pipeline.
 Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), an OM task can be considered as finding an alignment ($A$) that contains a collection of mappings.
@@
 However, while an alignment for OM only considers matched entities, OV tracks both matched and non-matched entities.
-Further, matched entities are composed of two subsets remain and update, and non-matched entities are composed of add and delete entities.
-In practice, we expect for OV that unchanged remain entities dominate.
-Figure 1: OM4OV pipeline.
-III-A Task Formulation
-The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or other relations, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$ [13].
-$\footnotesize A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$ (1)
+Furthermore, matched entities are composed of two subsets remain and update, while non-matched entities are composed of add and delete entities.
+In practice, we expect that unchanged remain entities dominate for OV.
+Figure 1: Overview of the OM4OV pipeline.
+Definition 3.1.
+The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as [6]: $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},s\leq c<1\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or other relations, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$.
 Similarly, an OV task can be formalised as finding two variants of an alignment $A_{match}$ and $A_{non-match}$ between ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$.
 While OM may have different mapping relations, OV narrows down the task and focuses only on the equivalence relation.
+Definition 3.2.
 We classify four subset alignments produced in the OV process, namely $A_{remain}$ ($A\odot$), $A_{update}$ ($A\otimes$), $A_{add}$ ($A\oplus$), and $A_{delete}$ ($A\ominus$).
-$A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c\geq s\}=A\odot\cup A\otimes$ (2)
-$A_{non-match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c<s\}=A\oplus\cup A\ominus$
-  1. 1.
+$A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},s\leq c\leq 1\}$ (1)
+$A_{non-match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},0\leq c<s\}$
 The remain and update entities are actually matched entities between different ontology versions.
 The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions.
-If the naming convention changes, we will classify ontology entities that have changed names according to systematic convention as update entities, by modelling their confidence $c$ as a high number more than $s$ but less than 1.
+If the naming convention changes, we will classify ontology entities that have changed names according to the systematic convention as update entities, by modelling their confidence $c$ as a high number more than $s$ but less than 1.
 The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions.
-Therefore, $A\odot$ and $A\otimes$ are defined as:
-    $A\odot=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c=1\}=A_{match}(c=1)$ (3)
-    $A\otimes=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},s\leq c<1\}=A_{match}(c\neq 1)$
-  2. 2.
+Definition 3.3.
+$A\odot$ and $A\otimes$ are defined as:
+$A\odot$ $=A_{match}(c=1)$ (2)
+$A\otimes$ $=A_{match}(s\leq c<1)$
 The add and delete entities are actually non-matched entities between different ontology versions.
@@
 Similarly, we can interpret the delete entities as non-matched entities in $O$.
-Therefore, $A\oplus$ and $A\ominus$ are defined as:
+Definition 3.4.
+$A\oplus$ and $A\ominus$ are defined as:
 $A\oplus=\{e2\in O^{\prime}$ $\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$ (3)
 $A\ominus=\{e1\in O$ $\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
-III-B Measurement
-Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision, recall, and F1 score.
-While precision measures matching correctness and recall measures matching completeness, there is an inherent trade-off between precision and recall.
-The F1 score offers a harmonic mean to balance the correctness and completeness.
-They are defined as:
-$Precison=\frac{|A\cap R|}{|A|}\quad Recall=\frac{|A\cap R|}{|R|}$ (5)
-$F_{1}\ Score=\frac{2}{Precison^{-1}+Recall^{-1}}$
-OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance.
+Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision (Prec), recall (Rec), and the F1 score (F1).
+While precision measures matching correctness and recall measures matching completeness, the F1 score offers a harmonic mean to balance correctness and completeness.
+Definition 3.5.
+OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance. They are defined as:
+$Prec\odot\otimes\oplus\ominus=\frac{|A\cap R|}{|A|}\quad Rec\odot\otimes\oplus\ominus=\frac{|A\cap R|}{|R|}\quad F_{1}\odot\otimes\oplus\ominus=\frac{2}{Prec^{-1}+Rec^{-1}}$ (4)
+4 Pipeline Evaluation
+4.1 Dataset Construction
+There is a dearth of lab-based datasets for the evaluation of OV.
+The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks; however, none of them includes benchmarks specifically designed for OV tasks.
+We propose an approach to constructing synthetic OV datasets from OM datasets.
+Figure 2 illustrates the generation of OM4OV datasets.
+Given a source ontology and target ontology with their reference alignment reference.xml, our approach is described in the following steps:
+Figure 2: Generation of OM4OV datasets.
+{bracketenumerate}
+The original OAEI datasets for OM provide two ontologies, $O_{s}$ and $O_{t}$. We choose one to be the intermediate ontology $O_{i}$. We retrieve all ontology entities from $O_{i}$.
+There are four possible changes in OV tasks: remain, update, add, and delete.
+Each ontology entity in $O_{i}$ is randomly assigned one of these.
+For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$.
+Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$.
+By construction, we also have the four sets as pairwise disjoint.
+For entities assigned to update, we need to generate the updated entity name.
+We should expect the new name to have a similar meaning to its original name.
+For example, the entity name “ConferenceVenuePlace” could be replaced with “Conference_hall” or “Conference_building”, but not the general names “Place” or “Location”.
+To achieve this goal, we retrieve all equivalent entities provided by reference.xml included in the original OAEI datasets and use them as a replacement synonym corpus.
+For entities whose names are unique identifiers or codes (and not textually meaningful names), we use their annotation properties (e.g.
+rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead.
+For entities that do not have synonyms in the generated list, we randomly re-assign them to remain, add, or delete.
+Based on the entity assignments, we generate four corresponding versioning references, namely vr-remain.xml, vr-update.xml, vr-add.xml, and vr-delete.xml.
+For entities assigned to remain, no operation is required.
+For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list.
+We generate $O$ and $O^{\prime}$ according to the following rules: {alphaenumerate}
+$O$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{add or }p\in\textit{add or }o\in\textit{add}\}$.
+That is, $O$ is constructed as $O_{i}$ without all the triples related to entities in the add list.
+Let $e$ be an entity and $A$ be a mapping of the form $\{e_{1}\rightarrow e^{\prime},e_{2}\rightarrow e^{\prime\prime},\ldots e_{n}\rightarrow e^{{}^{\prime\prime}...^{\prime}}\}$.
+Then $map(e,M)$ is defined to be $e^{\prime}$ if there is an $e^{\prime}$ such that $e\rightarrow e^{\prime}\in A$ and to be $e$ otherwise.
+Now, $O^{\prime}$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{delete}\textit{ or }p\in\textit{delete}\textit{ or }o\in\textit{delete}\}\cup\{(s^{\prime},p^{\prime},o^{\prime})|(s,p,o)\textit{ is a triple}\in O_{i}$ and $s^{\prime}=map(s,\textit{update})\textit{ or }p^{\prime}=map(p,\textit{update})\textit{ or }o^{\prime}=map(o,\textit{update})\}$.
+That is, $O^{\prime}$ is constructed as $O_{i}$ without all the triples related to entities in the delete list and updated for all the triples related to entities in the update list.
+We also incorporate the cross-reference mechanism into our OV testbed, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules: (1) $R_{or}$ is the original reference.xml that removes all the mappings related to add entities.
+(2) $R_{o^{\prime}r}$ is the original reference.xml that removes all the mappings related to delete entities and updates all the mappings related to update entities.
+Unlike the original OAEI datasets used for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed.
+This suits the dynamic nature of OV, where the changes vary between different versions.
+For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks.
+Table 1 lists the detailed information of the selected OAEI track used for the OV testbed.
+The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks.
+The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies.
+Figure 3 shows the number of entities in each track.
+Table 1: Selected OAEI tracks for the OV testbed.
+Track Domain Alignment Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 3
+Figure 3: Number of entities.
+4.2 System Implementation
+The OM4OV pipeline is implemented in Agent-OV, as an extension of Agent-OM [21].
+Agent-OM is an agent-powered LLM-based OM system.
+Its foundation framework is designed for traditional OM tasks.
+We extend Agent-OM with the OM4OV pipeline so that it can be used for OV tasks.
+Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [18] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [15], Alibaba Qwen [1], Google Gemma [7], and ChatGLM [25].
+We use the OpenAI model gpt-4o-mini (2024-07-18) [17] in our experiments.
+For performance using different LLMs, we refer the reader to [21], where we find that API-accessed LLMs generally perform better than open-source LLMs.
+The embedding chosen correspondingly is OpenAI text-embedding-ada-002 [8].
+4.3 Experimental Results
+Figure 4 shows the evaluation of Agent-OV and Agent-OM on the OV testbed.
+The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated.
+The results indicate that it is possible to unify the OM and OV tasks.
+Both Agent-OM and Agent-OV can produce an alignment with the precision, recall, and F1 score provided.
+However, the original OM systems require necessary modifications to fit into the OV tasks.
+We can see that Agent-OM always has an unusually high overall performance across all alignments.
+This is because the traditional OM system only captures matched entities between two ontologies, where this refers to the remain and update entities in the OV task.
+While remain entities commonly have a large proportion between two different versions, the measure of remain entities dominates the overall evaluation.
+We believe this is a common pitfall and causes skewed performance measurement when using the OM system in OV tasks.
+Agent-OV overcomes the skewed measure in Agent-OM.
+When decomposing the measure into four sub-measures, we can observe a more precise matching performance across different OV tasks.
+In general, we observe the matching performance is highest in remain, followed by add and delete, and relatively low in update.
+This trend is consistent across different tracks and ontologies.
+{bracketenumerate}
+The measures for remain are typically very close to 100%. This sub-measure is statistically significant and can be omitted in OV tasks.
+The measures for add and delete are generally good. Our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect non-matched entities.
+The measures for update show scope for improvement.
+This is because finding non-trivial alignments and appropriate similarity thresholds is not easy.
+Additionally, we observe a longer computation time in the Anatomy Track.
+Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still insufficient for some OV tasks.
+There is a need to optimise the pipeline for OV in large-scale ontologies.
+Figure 4: Evaluation of Agent-OM and Agent-OV on the OV testbed.
+Figure 5 shows the evaluation of using different similarity threshold methods in Agent-OV.
+We change the similarity threshold from 0.9 to 1 to evaluate its effect on matching performance.
+While OV entities are classified into four different categories, each category requires a corresponding sub-measurement.
+Similarly to OM tasks, changes in hyperparameter settings in OV tasks lead to a trade-off between precision and recall.
+Moreover, the hyperparameter settings also influence the sub-measures.
+We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities.
+Furthermore, detecting update entities and detecting add and delete operations are negatively correlated.
+Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities.
+Our explanations on these trends are as follows.
+Figure 5: Evaluation of using different similarity threshold methods in Agent-OV.
+{bracketenumerate}
 Within each sub-measure, the precision-recall trade-off still holds.
@@
 If we define $\Delta A$ as a universal change in OV, any changes in the alignments satisfy the following equation:
-$|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=-|\Delta A\oplus|=-|\Delta A\ominus|$ (7)
+$|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=|\Delta A\oplus|=|\Delta A\ominus|$ (6)
 We define the corresponding changes of $|A\cap R|$ in remain, update, add, and delete as $\Delta(A\cap R)\odot\otimes$ and $\Delta(A\cap R)\oplus\ominus$.
-  • For recall, there are direct relations between different sub-measures.
+For recall, there are direct relations between sub-measures.
 Recalldirect-$\odot$ and Recalltensor-$\otimes$ will increase with Recalldirect-$\oplus$ and Recallsymmetric-$\ominus$ decreasing and vice versa.
-    $Recall\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}$ (8)
-    $Recall\oplus\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|R|}$
-  • For precision, there are no direct relations between different sub-measures.
-The indirect relations, described qualitatively in the following as nondecreasing ($\uparrow$), nonincreasing ($\downarrow$), or either (?)?(?)( ?
+For precision, there are no direct relations between sub-measures.
+The indirect relations, described qualitatively in the following as increasing ($\uparrow$), decreasing ($\downarrow$), or either (?)?(?)( ?
 ), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
-    $Precision\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (9)
-    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\odot\otimes|\times|A|}<1\quad Precision\odot\otimes\uparrow$
-    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\odot\otimes|\times|A|}>1\quad Precision\odot\otimes\downarrow$
-    $Precision\oplus\ominus\ ?=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
-    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\oplus\ominus|\times|A|}>1\quad Precision\oplus\ominus\uparrow$
-    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\oplus\ominus|\times|A|}<1\quad Precision\oplus\ominus\downarrow$
-III-C Testbed
-There is a dearth of lab-based datasets for OV evaluation.
-The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none of them contain benchmarks designed for OV tasks.
-We propose an approach to constructing synthetic OV datasets from OM datasets.
-Fig.
-2 illustrates the generation of OM4OV datasets.
-Our approach is described in the following steps:
-Figure 2: Generation of OM4OV datasets.
-  1. 1.
-    The original OAEI datasets for OM provide two ontologies, $O_{s}$ and $O_{t}$. We choose one to be the intermediate ontology $O_{i}$. We retrieve all ontology entities from $O_{i}$.
-  2. 2.
-    There are four possible changes in OV tasks: remain, update, add, and delete.
-Each ontology entity in $O_{i}$ is randomly assigned one of these.
-For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$.
-Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$.
-By construction, we also have the four sets as pairwise disjoints.
-  3. 3.
-    For entities assigned to update, we need to generate the updated entity name.
-We should expect the new name to have a similar meaning to its original name.
-For example, the entity name “ConferenceVenuePlace” could be replaced with “Conference_hall” or “Conference_building”, but not the general names “Place” or “Location”.
-To achieve this goal, we retrieve all equivalent entities provided by reference.xml included in the original OAEI datasets and use them as a replacement synonym corpus.
-For those entities whose names are unique identifiers or codes (and not textually-meaningful names), we use their annotation properties (e.g.
-rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead.
-For those entities that do not have synonyms in the generated list, we randomly re-assign the entity to remain, add, or delete.
-  4. 4.
-    Based on the entity assignments, we generate four corresponding versioning references, namely vr-remain.xml, vr-update.xml, vr-add.xml, and vr-delete.xml.
-For entities assigned to remain, no operation is required.
-For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list.
-  5. 5.
-    We generate $O$ and $O^{\prime}$ according to the following two rules:
-    1. (a)
-      $O$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{add or }p\in\textit{add or }o\in\textit{add}\}$.
-That is, $O$ is constructed as $O_{i}$ without all the triples related to entities in the add list.
-    2. (b)
-      Let $e$ be an entity and $A$ be a mapping of the form $\{e_{1}\rightarrow e^{\prime},e_{2}\rightarrow e^{\prime\prime},\ldots e_{n}\rightarrow e^{{}^{\prime\prime}...^{\prime}}\}$.
-Then $map(e,M)$ is defined to be $e^{\prime}$ if there is an $e^{\prime}$ such that $e\rightarrow e^{\prime}\in A$ and to be $e$ otherwise.
-Now, $O^{\prime}$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{delete}\textit{ or }p\in\textit{delete}\textit{ or }o\in\textit{delete}\}\cup\{(s^{\prime},p^{\prime},o^{\prime})|(s,p,o)\textit{ is a triple}\in O_{i}$ and $s^{\prime}=map(s,\textit{update})\textit{ or }p^{\prime}=map(p,\textit{update})\textit{ or }o^{\prime}=map(o,\textit{update})\}$.
-That is, $O^{\prime}$ is constructed as $O_{i}$ without all the triples related to entities in the delete list and updated for all the triples related to entities in the update list.
-  6. 6.
-    We also incorporate the cross-reference mechanism into our OV testbed, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules: (1) $R_{or}$ is the original reference.xml that removes all the mappings related to add entities.
-(2) $R_{o^{\prime}r}$ is the original reference.xml that removes all the mappings related to delete entities and updates all the mappings related to update entities.
-Unlike the original OAEI datasets for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed.
-This suits the dynamic nature of OV, where the changes vary between different versions.
-For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks.
-The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks.
-Table I lists the detailed information of the selected OAEI track used for the OV testbed.
-The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies.
-TABLE I: Selected OAEI track for the OV testbed.
-Track Domain Number of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 3
-III-D Cross-Reference Mechanism
-Often, ontology creators provide cross-references to other ontologies to enhance interoperability.
-For example, the cross-reference between the CMT ontology and the Conference ontology is provided along with the CMT ontology.
-Reusing these cross-references developed for OM tasks, we propose a novel mechanism to reduce matching candidates and also to improve overall OV performance.
-Fig.
-3 illustrates the cross-reference mechanism used in the OM4OV pipeline.
+$Rec\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}\quad Prec\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (7)
+$Rec\oplus\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|R|}\quad Prec\oplus\ominus\ ?=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
+5 Pipeline Optimisation
+5.1 Motivation
+Often, ontology creators provide cross-references to other ontologies to enhance interoperability and facilitate integration.
+For example, a cross-reference between the CMT ontology and the Conference ontology is provided alongside the CMT ontology.
+Reusing the cross-references developed for OM tasks, we propose a novel mechanism to reduce the number of matching candidates and also to improve overall OV performance.
+5.2 Approach
+Figure 6 illustrates the cross-reference mechanism used in the OM4OV pipeline.
 We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$.
@@
 $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$).
-The prior matching will be part of the final alignment, while the known non-matching will be completely removed in the subsequent OV process.
+The prior matching will be incorporated into the final alignment, while the known non-matching regions will be removed entirely in the subsequent OV process.
 The OV process then only determines the posterior alignment.
 In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities.
-Matching performance is also improved by using these known mappings.
+Matching performance is also enhanced by utilising these established mappings.
 Since prior matches are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain.
-On the other hand, while the known non-matches are removed from the OV process, this reduces the complexity of detecting the posterior alignment.
+On the other hand, while the known non-matches are removed from the OV process, this simplifies the detection of the posterior alignment.
 Figure 6: Cross-reference mechanism.
+Definition 5.1.
 Given a reference ontology ($O_{r}$) with an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), two cross-references between $O$ and $O_{r}$ ($R_{or}$) and between $O^{\prime}$ and $O_{r}$ ($R_{o^{\prime}r}$) are defined as:
-$R_{or}$ =$=\{(e1,e3,\equiv,c)|e1,e3\in O_{r}\cap O,c\geq s\}$ (10)
-$R_{o^{\prime}r}$ =$=\{(e2,e3,\equiv,c)|e2,e3\in O_{r}\cap O^{\prime},c\geq s\}$
+$R_{or}$ =$=\{(e1,e3,\equiv,c)|e1,e3\in O_{r}\cap O,s\leq c\leq 1\}$ (8)
+$R_{o^{\prime}r}$ =$=\{(e2,e3,\equiv,c)|e2,e3\in O_{r}\cap O^{\prime},s\leq c\leq 1\}$
+Definition 5.2.
 We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV.
@@
 After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$).
-Therefore, $A_{match}$ in OV can be decomposed into two parts:
+In this setting, $A_{match}$ in OV can be decomposed into two parts: $A_{\pi}$ and $A_{\pi^{*}}$.
 $A_{match}=A_{\pi}\cup A_{\pi^{*}}$ (9)
-  1. 1.
+Definition 5.2.1.
 $A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$.
 The equivalence relation is transitive, so for $e1\in O$, $e2\in O^{\prime}$, and $e3\in O_{r}$, if $e1=e3$ and $e2=e3$ then $e1=e2$.
-    $A_{\pi}=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},c\geq s\}$ (12)
-  2. 2.
 $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference.
@@
 As a result, $A_{\pi^{*}}$ can be defined within a smaller scope:
-    $A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},c\geq s\}$ (13)
+$A_{\pi}$ $=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},s\leq c\leq 1\}$ (10)
+$A_{\pi^{*}}$ =$=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},s\leq c\leq 1\}$
+Definition 5.2.2.
 […]
 $A_{\pi}$ $=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup...\cup(R_{orn}\cap R_{o^{\prime}rn})$ (11)
-$A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}$
-$\text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup...O_{rn}$
-IV Implementation & Evaluation
-The OM4OV pipeline is implemented in Agent-OV, a variant of Agent-OM [14].
-Agent-OM is an agent-powered LLM-based OM system.
-Its foundation framework is designed for traditional OM tasks.
-We extend the original framework with the OM4OV pipeline so that it can be used to handle OV tasks.
-As an extension of Agent-OM, Agent-OV also supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [15], Anthropic Claude [16], and Mistral AI [17], as well as open-source LLMs Meta Llama [18], Google Gemma [19], and Alibaba Qwen [20].
-For the performance of Agent-OV using different LLMs, we refer the reader to [14], where we find that API-accessed LLMs generally perform better than open-source LLMs.
-IV-A Evaluation of OV testbed
-Fig.
-4 shows the evaluation of the Agent-OV system on the OV testbed.
-We use the OpenAI model gpt-4o-mini (2024-07-18) [21] for evaluation.
-The embedding chosen correspondingly is OpenAI text-embedding-ada-002 [22].
-The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated.
-We run the OV testbed for 5 iterations, simulating 5 distinct versions for each ontology.
-The results indicate that it is possible to unify the OM and OV tasks.
-With the necessary modifications, the OM systems can also be used in OV tasks.
-(a) Anatomy Track: human
-(b) Anatomy Track: mouse
-(c) Conference Track: cmt
-(d) Conference Track: conference
-(e) Conference Track: confof
-(f) Conference Track: edas
-(g) Conference Track: ekaw
-(h) Conference Track: iasted
-(i) Conference Track: sigkdd
-(j) MSE Track: EMMO
-(k) MSE Track: MatInfo
-(l) MSE Track: MatOnto
-Figure 4: Evaluation of OV testbed.
-  1. 1.
-    Our experiments show that the OM system achieves acceptable performance in tracking changes over different versions of the ontologies.
-Interestingly, the performance of our system on OV tasks is generally better than on OM tasks.
-This could be due to the fact that the ontologies used in the OV tasks are two versions of the same ontology that share consistent ontology design patterns.
-  2. 2.
-    For the four micro-level sub-measures, we observe:
-    1. (a)
-      The performance is highest in remain, followed by add and delete, and it is relatively low in update. This trend is consistent across different tracks and ontologies.
-    2. (b)
-      The measurements for remain are commonly very close to 100% and not statistically significant in OV tasks.
-    3. (c)
-      The measures for add and delete are generally good. This may be because our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect missing mappings.
-    4. (d)
-      The measurements for update show scope for improvement.
-This is because finding non-trivial alignments and appropriate similarity thresholds is not easy.
-Unlike for OM tasks, we believe mixing remain and update in OV tasks is a common pitfall and can cause skewed performance measurement.
-  3. 3.
-    For the computational time, we observe a longer computation time for the large-scale ontologies in the Anatomy Track.
-Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still not sufficient for some OV tasks.
-IV-B Evaluation of cross-reference mechanism
-Fig.
-5 shows the comparison of the OV performance with and without using the cross-reference (CR) mechanism on the same OV testbed we analysed in Section IV-A.
-We can see that the CR mechanism significantly improves the capability of Agent-OV to detect update entities and slightly improves the performance on detecting add and delete entities, with no effect on the remain entities.
-More specifically, the performance improvement is mainly caused by the recall for update entities, while this is more evident in the precision for add and delete entities.
-(a) Remain-Precision
-(b) Update-Precision
-(c) Add-Precision
-(d) Delete-Precision
-(e) Remain-Recall
-(f) Update-Recall
-(g) Add-Recall
-(h) Delete-Recall
-(i) Remain-F1 Score
-(j) Update-F1 Score
-(k) Add-F1 Score
-(l) Delete-F1 Score
-Figure 5: Evaluation of cross-reference (CR) mechanism.
-IV-C Evaluation of hyperparameter settings
-Fig.
-6 shows the comparison of using different similarity thresholds in the CMT Ontology versioning.
-Similar to OM tasks, changes in hyperparameter settings in OV tasks will lead to a trade-off in precision and recall.
-Moreover, the hyperparameter settings also influence the sub-measures.
-We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities.
-Also, detecting update entities and detecting add and delete is negatively correlated.
-Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities.
-The result of the experiment also shows that our cross-reference mechanism is less sensitive to the similarity threshold, which means that it can be useful in scenarios where the optimal similarity threshold is unclear or hard to determine.
-Figure 6: Evaluation of different similarity thresholds.
-V Discussion
+$A_{\pi^{*}}$ =$=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}\ \text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup...O_{rn}$
+5.3 Evaluation
+Figure 7 shows the comparison of the OV performance with and without the cross-reference (CR) mechanism on the same OV testbed that we analysed in Section 4.
+We can see that the CR mechanism significantly enhances the capability of Agent-OV to detect update entities and slightly improves performance in detecting add and delete entities, with no effect on remain entities.
+More specifically, performance improvement is mainly attributed to recall for update entities, while this is more evident in the precision for add and delete entities.
+Figure 7: Evaluation of the cross-reference (CR) mechanism on the OV testbed.
+Figure 8 shows the comparison of using different similarity thresholds in the CMT Ontology versioning.
+The results of the experiment also show that our cross-reference mechanism is less sensitive to the similarity threshold, which means it can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
+Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds.
+6 Discussion
 So far, we have experimentally validated the OM4OV pipeline with a novel cross-reference mechanism for pipeline optimisation.
@@
 Are these mappings actually false?
-  1. 1.
-    False mappings in OV can be a wrong ontology design choice.
+{bracketenumerate}
+False mappings in OV can result from a flawed ontology design choice.
 In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset.
@@
 cmt:writtenBy => cmt:isWrittenBy
-  2. 2.
-    False mappings in OV can be an ambiguous “equivalent” relationship.
+False mappings in OV can create an ambiguous “equivalent” relationship.
 In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference.
 However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair.
-This interpretation is not wrong but follows a different ontology design pattern.
-It is important to notice that the term ”equivalent” in OV is weaker than that in OM.
+This interpretation is not wrong, but follows a different ontology design pattern.
+It is vital to notice that the term “equivalent” in OV is weaker than that in OM.
 OV allows for roughly “equivalent” mappings.
@@
 cmt:ConferenceChair => cmt:General_Chair
-  3. 3.
-    False mappings in OV can be an inappropriate setup on the similarity threshold.
+False mappings in OV can be an inappropriate setting of the similarity threshold.
 In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$.
@@
 s = 0.90, cmt:SubjectArea => cmt:Topic
-VI Limitations
+7 Limitations
+{bracketenumerate}
 We focus mainly on tracking conceptual changes in classes and properties (e.g.
-adding, deleting, or updating).
+adding, deleting, or updating class and properties).
 In practice, there are also internal relationship changes (e.g.
@@
 For example, using a harmonic mean to combine sub-measures.
-VII Conclusion
+8 Conclusion
 In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with the necessary modifications.
@@
 The new pipeline (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the matching candidates, and (3) improves overall performance.
-We incorporate the OM4OV pipeline into a new OV system called Agent-OV, a functional extension of Agent-OM to handle OV tasks.
-Evaluations of three OAEI datasets validate the feasibility and reliability of our system.
-We also argue that the false mappings detected by OV systems are not necessarily actual false mappings.
+We incorporate the OM4OV pipeline into a new OV system called Agent-OV, which serves as a functional extension of Agent-OM to handle OV tasks.
+Evaluations of our OAEI tested validate the feasibility and reliability of our system.
+Finally, we argue that the false mappings detected by OV systems are not necessarily actual false mappings.
 Our approach is compatible with ontologies using different versioning methods (using URIs or additional versioning triples), and even ontologies missing or without version information.
 Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies.
-In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time (e.g.
-smart building ontologies Brick Schema [23], RealEstateCore [24], and ASHRAE Standard 223P [25]).
-Supplemental Material Statement
-Agent-OV is an extended version of Agent-OM.
-Agent-OM can be downloaded from https://github.com/qzc438-research/ontology-llm.
-The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01).
-The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
-Acknowledgment
-The authors would like to thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper.
-The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
+In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time.
+For example, smart building ontologies Brick Schema [3], RealEstateCore [11], and ASHRAE Standard 223P [16].

### v4

@@
 Despite the long-standing recognition of ontology versioning (OV) as a crucial component for efficient ontology management, the growing size of ontologies and accumulating errors caused by manual labour overwhelm current OV approaches.
-In this paper, we propose yet another approach to performing OV using existing ontology matching (OM) techniques and systems.
+In this paper, we propose a fresh approach to performing OV using existing ontology matching (OM) techniques and systems.
 We introduce a unified OM4OV pipeline.
-From an OM perspective, we reconstruct a new task formulation and measurement for OV tasks.
+From an OM perspective, we reconstruct a new task formulation and measurements for OV tasks.
 Building upon the prior alignment(s) from OM, we propose a pipeline optimisation method called the cross-reference (CR) mechanism to enhance overall OV performance.
-We experimentally validate the OM4OV pipeline and the cross-reference mechanism in the OV tested originating from the Ontology Alignment Evaluation Initiative (OAEI) datasets.
-We also discuss insights into OM used for OV tasks, where some false mappings detected by OV systems are not actually untrue.
+We experimentally validate the OM4OV pipeline and the cross-reference mechanism in an OV testbed originating from the Ontology Alignment Evaluation Initiative (OAEI) datasets.
+We also discuss insights into OM used for OV tasks, where some apparent false mappings detected by OV systems are not actually untrue.
 keywords:
@@
 Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications [8].
-An ontology is not static, and the need for version control arises with its birth.
+An ontology is not static, and the need for version control arises with its origination.
 While web data is dynamic, any ontology used must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation [13].
@@
 Ontology versioning (OV) aims to distinguish and recognise changes between different ontology versions.
-By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology, can apply the correct changes correspondingly [14].
+By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology can apply the correct changes correspondingly [13].
 While various methods for OV have been developed, one approach is to extend the ontology itself with internal version information.
@@
 IRI) or a specific version number (e.g.
-owl:versionInfo) to be distinguished from other versions; each ontology entity may have a new triple to record its current status (e.g.
+owl:versionInfo) to be distinguished from other versions; each ontology entity may have a new annotation to record its current status (e.g.
 owl:DeprecatedClass and owl:DeprecatedProperty); or every triple could be extended with a 4th dimension populated with a triple timestamp, similar to the RDF-star schema.
 Alternatively, version information can be recorded in change logs.
-Change logs can take the form of notebooks, an extensible markup language (XML), or a knowledge graph (KG).
+Change logs can take the form of free-text notebooks, an extensible markup language (XML) document, or a knowledge graph (KG).
 However, maintaining version information in the ontology can be time-consuming and labour-intensive.
@@
 A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts.
-Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or has a separate change log.
+Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or that it has a separate change log.
 In such cases, current approaches have limited capabilities in detecting incorrect version information or identifying missing version information.
@@
 However, OM techniques and systems cannot be directly applied to OV tasks because OM and OV have some significant differences.
-(1) The OM input is two distinct ontologies, while the OV input is two different versions of a single ontology.
-The output of OM is a set of entity mappings, whereas the output of OV is two sets: changed entities and unchanged entities.
-(2) OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity.
-OM determines matched entities between two different ontologies, while OV distinguishes add, delete, remain, and update entities between different versions in one ontology.
-To address these challenges, our goal is to investigate the complementary relationship between these two tasks and to develop a unified approach to shifting OV tasks towards OM tasks.
+The OM input is two distinct ontologies, while the OV input is expected to be two different versions of a single ontology.
+The output of OM is a set of entity mappings, whereas the output of OV comprises two sets: changed entities and unchanged entities.
+OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity.
+OM determines matched entities between two different ontologies, while OV distinguishes add, delete, remain, and update entities between different versions of one ontology.
+To address these challenges, our goal is to investigate the complementary relationship between these two tasks and to develop a unified approach to shifting OV tasks towards OM tasks, thereby leveraging the deep research legacy of OM to assist with OV.
 Specifically, our key contributions include:
@@
   • We implement the OM4OV pipeline and the cross-reference mechanism in a proof-of-concept system and experimentally evaluate its performance.
-  • We find that some false mappings detected by OV systems are not actually untrue, but can be caused by a flawed choice of ontology design, an ambiguous “equivalent” relationship, or an inappropriate setting of the similarity threshold.
+  • We argue that some false mappings detected by OV systems are not actually untrue, but can be caused by a flawed choice of ontology design, an ambiguous “equivalent” relationship, or an inappropriate setting of a similarity threshold.
 The remainder of the paper is organised as follows.
@@
 Section 3 introduces our novel OM4OV pipeline, while Section 4 evaluates the proposed pipeline.
-We propose our pipeline optimisation method (i.e.
-cross-reference mechanism) and experimentally evaluate its performance in Section 5.
+We propose the cross-reference mechanism as a pipeline optimisation method and experimentally evaluate its performance in Section 5.
 Section 6 and 7 discuss our implications and current limitations.
@@
 These approaches have largely been adopted by the later OWL Web Ontology Language [4], where a set of annotation properties related to version information is defined.
-These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify the entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare the deprecated entities.
+These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify an entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare deprecated entities.
 Later, the ontology language $\tau$OWL [22] was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology.
@@
 To the best of our knowledge, our work is the first to systematically analyse and utilise OM for OV.
+The output of our novel automatic OV detection could be easily recorded by any of the current approaches presented above.
 3 OM4OV Pipeline
@@
 Figure 1: Overview of the OM4OV pipeline.
-Definition 3.1.
 The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as [6]: $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},s\leq c<1\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or other relations, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$.
@@
 Definition 3.1.
-We classify four subset alignments produced in the OV process, namely $A_{remain}$ ($A\odot$), $A_{update}$ ($A\otimes$), $A_{add}$ ($A\oplus$), and $A_{delete}$ ($A\ominus$).
+An OV task is defined as:
 $A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},s\leq c\leq 1\}$ (1)
 $A_{non-match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},0\leq c<s\}$
+We classify four subset alignments produced in the OV process, namely $A_{remain}$ ($A\odot$), $A_{update}$ ($A\otimes$), $A_{add}$ ($A\oplus$), and $A_{delete}$ ($A\ominus$).
 The remain and update entities are actually matched entities between different ontology versions.
@@
 The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions.
-Definition 3.3.
-$A\odot$ and $A\otimes$ are defined as:
-$A\odot$ $=A_{match}(c=1)$ (2)
+Definition 3.1.1.
+$A_{match}$, $A\odot$ and $A\otimes$ are defined as:
+$A_{match}=A\odot\cup A\otimes$ (2)
+$A\odot=A_{match}(c=1)$
 $A\otimes=A_{match}(s\leq c<1)$
@@
 Similarly, we can interpret the delete entities as non-matched entities in $O$.
-Definition 3.4.
-$A\oplus$ and $A\ominus$ are defined as:
-$A\oplus=\{e2\in O^{\prime}$ $\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$ (3)
-$A\ominus=\{e1\in O$ $\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
+Definition 3.1.2.
+$A_{non-match}$, $A\oplus$, and $A\ominus$ are defined as:
+$A_{non-match}=A\oplus\cup A\ominus$ (3)
+$A\oplus=\{e2\in O^{\prime}\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$
+$A\ominus=\{e1\in O\phantom{{}^{\prime}}\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
 Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision (Prec), recall (Rec), and the F1 score (F1).
 While precision measures matching correctness and recall measures matching completeness, the F1 score offers a harmonic mean to balance correctness and completeness.
+OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance.
 Definition 3.2.
-OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance. They are defined as:
-$Prec\odot\otimes\oplus\ominus=\frac{|A\cap R|}{|A|}\quad Rec\odot\otimes\oplus\ominus=\frac{|A\cap R|}{|R|}\quad F_{1}\odot\otimes\oplus\ominus=\frac{2}{Prec^{-1}+Rec^{-1}}$ (4)
+The measures of OV task are defined as:
+$Prec\odot|\otimes|\oplus|\ominus=\frac{|A\cap R|}{|A|}\quad Rec\odot|\otimes|\oplus|\ominus=\frac{|A\cap R|}{|R|}\quad F_{1}\odot|\otimes|\oplus|\ominus=\frac{2}{Prec^{-1}+Rec^{-1}}$ (4)
 4 Pipeline Evaluation
 4.1 Dataset Construction
-There is a dearth of lab-based datasets for the evaluation of OV.
-The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks; however, none of them includes benchmarks specifically designed for OV tasks.
+There is a dearth of benchmark datasets for the evaluation of OV.
+The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks; but none are specifically designed for OV tasks.
 We propose an approach to constructing synthetic OV datasets from OM datasets.
 Figure 2 illustrates the generation of OM4OV datasets.
-Given a source ontology and target ontology with their reference alignment reference.xml, our approach is described in the following steps:
+Given a source ontology and a target ontology with their reference alignment reference.xml, our approach is described in the following steps:
 Figure 2: Generation of OM4OV datasets.
 {bracketenumerate}
-The original OAEI datasets for OM provide two ontologies, $O_{s}$ and $O_{t}$. We choose one to be the intermediate ontology $O_{i}$. We retrieve all ontology entities from $O_{i}$.
-There are four possible changes in OV tasks: remain, update, add, and delete.
-Each ontology entity in $O_{i}$ is randomly assigned one of these.
+The original OAEI datasets for OM provide two ontologies, the source $O_{s}$ and the target $O_{t}$.
+We choose one to be the intermediate ontology $O_{i}$.
+We retrieve all ontology entities from $O_{i}$.
+There are four possible entity changes in OV tasks: remain, update, add, and delete.
+Each ontology entity in $O_{i}$ is randomly assigned to one of these.
 For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$.
@@
 By construction, we also have the four sets as pairwise disjoint.
-For entities assigned to update, we need to generate the updated entity name.
+For entities assigned to update, we need to generate a new name for the updated entity.
 We should expect the new name to have a similar meaning to its original name.
@@
 For entities whose names are unique identifiers or codes (and not textually meaningful names), we use their annotation properties (e.g.
-rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead.
-For entities that do not have synonyms in the generated list, we randomly re-assign them to remain, add, or delete.
+rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead of the code.
+For entities in update that do not have synonyms in the corpus, we randomly re-assign them to remain, add, or delete.
 Based on the entity assignments, we generate four corresponding versioning references, namely vr-remain.xml, vr-update.xml, vr-add.xml, and vr-delete.xml.
+For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list.
 For entities assigned to remain, no operation is required.
-For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list.
 We generate $O$ and $O^{\prime}$ according to the following rules: {alphaenumerate}
@@
 That is, $O$ is constructed as $O_{i}$ without all the triples related to entities in the add list.
-Let $e$ be an entity and $A$ be a mapping of the form $\{e_{1}\rightarrow e^{\prime},e_{2}\rightarrow e^{\prime\prime},\ldots e_{n}\rightarrow e^{{}^{\prime\prime}...^{\prime}}\}$.
-Then $map(e,M)$ is defined to be $e^{\prime}$ if there is an $e^{\prime}$ such that $e\rightarrow e^{\prime}\in A$ and to be $e$ otherwise.
+Let $e$ be an entity and $M$ be a mapping of the form $\{e_{1}\rightarrow e^{\prime},e_{2}\rightarrow e^{\prime\prime},\ldots e_{n}\rightarrow e^{{}^{\prime\prime}...^{\prime}}\}$.
+Then $map(e,M)$ is defined to be $e^{\prime}$ if there is an $e^{\prime}$ such that $e\rightarrow e^{\prime}\in M$ and to be $e$ otherwise.
 Now, $O^{\prime}$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{delete}\textit{ or }p\in\textit{delete}\textit{ or }o\in\textit{delete}\}\cup\{(s^{\prime},p^{\prime},o^{\prime})|(s,p,o)\textit{ is a triple}\in O_{i}$ and $s^{\prime}=map(s,\textit{update})\textit{ or }p^{\prime}=map(p,\textit{update})\textit{ or }o^{\prime}=map(o,\textit{update})\}$.
@@
 For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks.
+For reproducibility, we use a fixed seed to control the randomness in steps (3) and (4).
+4.2 Experiment Setup
 Table 1 lists the detailed information of the selected OAEI track used for the OV testbed.
@@
 Figure 3 shows the number of entities in each track.
+The fixed seed for the OV testbed is set to 42.
 Table 1: Selected OAEI tracks for the OV testbed.
@@
 Figure 3: Number of entities.
-4.2 System Implementation
 The OM4OV pipeline is implemented in Agent-OV, as an extension of Agent-OM [19].
@@
 Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [16] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [14], Alibaba Qwen [1], Google Gemma [7], and ChatGLM [23].
-We use the OpenAI model gpt-4o-mini (2024-07-18) [17] in our experiments.
 For performance using different LLMs, we refer the reader to [19], where we find that API-accessed LLMs generally perform better than open-source LLMs.
-The embedding chosen correspondingly is OpenAI text-embedding-ada-002 [8].
-4.3 Experimental Results
-Figure 4 shows the evaluation of Agent-OV and Agent-OM on the OV testbed.
+We choose to use the Meta open-source model llama-3-8b for our experiments in this paper, eschewing the higher-end models which offer small improvements that may not justify the financial investment.
+We use embeddings from the same LLM model to further avoid expensive API calls and thereby provide an entirely free-to-use version of Agent-OV.
+4.3 Results
+Figure 4 evaluates Agent-OV and Agent-OM on the OV testbed.
 The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated.
-The results indicate that it is possible to unify the OM and OV tasks.
-Both Agent-OM and Agent-OV can produce an alignment with the precision, recall, and F1 score provided.
-However, the original OM systems require necessary modifications to fit into the OV tasks.
+The results indicate that it is possible to unify the OM and OV tasks as developed in this paper.
+Both Agent-OM and Agent-OV can produce an alignment with the precision, recall, and F1 scores of Figure 4.
+However, the original OM systems require necessary modifications to tackle OV tasks.
 We can see that Agent-OM always has an unusually high overall performance across all alignments.
 This is because the traditional OM system only captures matched entities between two ontologies, where this refers to the remain and update entities in the OV task.
-While remain entities commonly have a large proportion between two different versions, the measure of remain entities dominates the overall evaluation.
-We believe this is a common pitfall and causes skewed performance measurement when using the OM system in OV tasks.
+While remain entities commonly comprise the major proportion of entities over two different versions, the measure of remain entities dominates the overall evaluation.
+We believe this is a common situation and causes misleading performance measurement when using the OM system in OV tasks.
 Agent-OV overcomes the skewed measure in Agent-OM.
@@
 Figure 4: Evaluation of Agent-OM and Agent-OV on the OV testbed.
-Figure 5 shows the evaluation of using different similarity threshold methods in Agent-OV.
+Figure 5 studies the effect of varying similarity thresholds in Agent-OV.
 We change the similarity threshold from 0.9 to 1 to evaluate its effect on matching performance.
@@
 Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities.
-Our explanations on these trends are as follows.
-Figure 5: Evaluation of using different similarity threshold methods in Agent-OV.
+Our explanations for these trends are as follows.
+Figure 5: Evaluation varying similarity thresholds in Agent-OV. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
 {bracketenumerate}
@@
 For precision, there are no direct relations between sub-measures.
-The indirect relations, described qualitatively in the following as increasing ($\uparrow$), decreasing ($\downarrow$), or either (?)?(?)( ?
-), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
-$Rec\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}\quad Prec\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (7)
-$Rec\oplus\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|R|}\quad Prec\oplus\ominus\ ?=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
+The indirect relations, described qualitatively in the following as increasing ($\uparrow$), decreasing ($\downarrow$), or uncertain ($\sim$), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
+$Rec\odot|\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot|\otimes|\uparrow}{|R|}\quad Prec\odot|\otimes\sim=\frac{|A\cap R|+|\Delta(A\cap R)\odot|\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (7)
+$Rec\oplus|\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus|\ominus|\downarrow}{|R|}\quad Prec\oplus|\ominus\sim=\frac{|A\cap R|-|\Delta(A\cap R)\oplus|\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
 5 Pipeline Optimisation
@@
 5.3 Evaluation
-Figure 7 shows the comparison of the OV performance with and without the cross-reference (CR) mechanism on the same OV testbed that we analysed in Section 4.
+Figure 7 compares the OV performance with and without the cross-reference (CR) mechanism on the same OV testbed that we analysed in Section 4.
 We can see that the CR mechanism significantly enhances the capability of Agent-OV to detect update entities and slightly improves performance in detecting add and delete entities, with no effect on remain entities.
 More specifically, performance improvement is mainly attributed to recall for update entities, while this is more evident in the precision for add and delete entities.
-Figure 7: Evaluation of the cross-reference (CR) mechanism on the OV testbed.
-Figure 8 shows the comparison of using different similarity thresholds in the CMT Ontology versioning.
-The results of the experiment also show that our cross-reference mechanism is less sensitive to the similarity threshold, which means it can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
+Figure 7: Precision, Recall and F1 evaluation of the cross-reference (CR) mechanism on the OV testbed.
+Figure 8 compares the effects of different similarity thresholds in the CMT Ontology versioning.
+The results of the experiment show that our cross-reference mechanism is less sensitive to the similarity threshold than the original OM4OV pipeline, which means the cross-reference mechanism can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
 Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds.
+We apply a 1-D Gaussian filter for smoothing.
+The standard deviation for a Gaussian kernel is 1 (i.e.
+$sigma=1$).
 6 Discussion
@@
 In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset.
-This issue is caused by incorrect name conventions.
+This issue is caused by inconsistent naming convention practice.
 In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer).
-However, using cmt:writtenBy for paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author).
+However, using cmt:writtenBy for the paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author).
 Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other.
@@
 OV allows for roughly “equivalent” mappings.
-The entities mapped in OV can slightly alter their meanings in response to changes in the domain knowledge understanding.
+The entities mapped in OV can slightly alter their meanings in response to changes in domain understanding or evolution of natural language in the domain.
 cmt:Chairman => cmt:Chair
 cmt:ConferenceChair => cmt:General_Chair
-False mappings in OV can be an inappropriate setting of the similarity threshold.
+False mappings in OV can arise from an inappropriate setting of the similarity threshold.
 In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$.
 Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent.
-For example, the similarity threshold is relatively higher in the biomedical domain to ensure each terminology is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
+For example, the similarity threshold could be relatively higher in the biomedical domain to ensure each term is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
 s = 0.95, cmt:SubjectArea => None
@@
 7 Limitations
-{bracketenumerate}
 We focus mainly on tracking conceptual changes in classes and properties (e.g.
@@
 8 Conclusion
-In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with the necessary modifications.
+In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with minor modifications.
 We propose a novel OM4OV pipeline with a cross-reference mechanism that leverages OM for OV.
@@
 We incorporate the OM4OV pipeline into a new OV system called Agent-OV, which serves as a functional extension of Agent-OM to handle OV tasks.
-Evaluations of our OAEI tested validate the feasibility and reliability of our system.
+Evaluation experiments on our OAEI testbed validate the feasibility and reliability of our system.
 Finally, we argue that the false mappings detected by OV systems are not necessarily actual false mappings.
-Our approach is compatible with ontologies using different versioning methods (using URIs or additional versioning triples), and even ontologies missing or without version information.
+Our approach is compatible with ontologies using different versioning methods, that is, using URIs or additional versioning triples, and ontologies missing or without version information.
 Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies.